### v1

# Michel F. C. Haddad

Abstract
We propose a difference-in-differences (DiD) method for a time-varying continuous treatment and multiple time periods. Our framework assesses the average treatment effect on the treated (ATET) when comparing two non-zero treatment doses. The identification is based on a conditional parallel trend assumption imposed on the mean potential outcome under the lower dose, given observed covariates and past treatment histories. We employ kernel-based ATET estimators for repeated cross-sections and panel data adopting the double/debiased machine learning framework to control for covariates and past treatment histories in a data-adaptive manner. We also demonstrate the asymptotic normality of our estimation approach under specific regularity conditions. In a simulation study, we find a compelling finite sample performance of undersmoothed versions of our estimators in setups with several thousand observations.
Keywords: Causal inference, Continuous treatment, Difference-in-differences, Machine learning, Parallel trends.
1 Introduction
Difference-in-differences (DiD) is a cornerstone method for causal inference, i.e., the evaluation of the impact of a treatment (such as vaccination) on an outcome of interest (such as mortality), in observational studies, provided that outcomes may be observed both before and after the treatment introduction. In such studies, the observed outcomes of treated individuals before and after a treatment typically do not allow researchers to directly infer the treatment effect. This is due to confounding time trends in the treated individuals’ counterfactual outcomes that would have occurred without the treatment. The DiD approach tackles this issue based on a control group not receiving the treatment and the so-called parallel trends assumption, imposing that the treated group’s unobserved counterfactual outcomes under non-treatment follow the same observed mean outcome trend of the control group. While the canonical DiD setup considers a binary treatment definition (treatment versus no treatment), in many empirical applications, treatments are continuously distributed - e.g., the vaccination rate in a region. Furthermore, the treatment intensity might vary over time and a strict control group with a zero treatment might not be available across all or even any treatment period, as also argued in de Chaisemartin et al. (2024), who point to taxes, tariffs, or prices as possible treatment variables with strictly non-zero doses.
To address these limitations, we propose an DiD approach for a time-varying continuous treatment across multiple time periods, enabling the assessment of the average treatment effect on the treated (ATET) when comparing two non-zero treatment doses, such as a higher versus a lower vaccination rate in a region. Identification relies on a conditional parallel trend assumption imposed on the mean potential outcome under the lower dose, given observed covariates and past treatment histories. This assumption requires that, conditional on covariates and previous treatments, the counterfactual outcomes of the treatment group (receiving the higher treatment dose) under the lower dose follow the same mean outcome trend as in the control group (receiving the lower treatment dose). We note that this parallel trend assumption on counterfactual outcomes under a positive treatment dose is more stringent than that on non-treated outcomes, see e.g., the discussion in Fricke (2017), which is the trade-off for DiD-based identification of causal effects based on comparisons of non-zero treatment doses.
We suggest novel semiparametric kernel-based DiD estimators for both repeated cross-sections, in which subjects differ across time periods; and panel data, in which the same subjects are repeatedly observed across periods. The estimators adopt the double/debiased machine learning (DML) framework introduced in Chernozhukov et al. (2018) to control for covariates and past treatment histories in a data-adaptive manner, which appears attractive in applications where the number of potential control variables is large relative to the sample size. Thus, we estimate the models for the conditional mean outcomes and conditional treatment densities - also known as generalized treatment propensity scores, based on machine learning and use these nuisance parameters as plug-ins in doubly robust (DR) score functions (see for instance Robins et al. (1994) and Robins and Rotnitzky (1995), which are tailored to DiD-based ATET estimation).
These DR scores include kernel functions on the continuous treatment as also considered in Kennedy et al. (2017) and Colangelo and Lee (2023), when imposing a selection-on-observables-assumption (rather than a parallel trend assumption) for identification. Under an appropriately shrinking bandwidth as the sample size increases, we demonstrate that our estimators satisfy Neyman-orthogonality, implying that they are substantially robust to approximation errors in the nuisance parameters under specific regularity conditions - e.g., approximate sparsity when using lasso regression for nuisance parameter estimation as considered in Belloni et al. (2014)). Following Chernozhukov et al. (2018), we also apply cross-fitting such that the nuisance parameters and scores are not simultaneously estimated in the same parts of the data, to avoid overfitting. To account for the biases introduced by the kernel functions, we use undersmoothed bandwidths, implying that the biases vanish asymptotically. We show that under particular regularity conditions, the DML kernel estimators satisfy asymptotic normality, which also provides us with an asymptotic formula for variance estimation. We also conduct a simulation study, which points to a compelling finite sample behavior of undersmoothed versions of our estimators under a sample size of several thousand observations.
Our study adds to a growing literature on the semi- and nonparametric DiD-based identification of causal effects of continuously distributed treatments. This framework avoids potential misspecification errors of classical linear two-way fixed effects models - see, for instance, the discussion in De Chaisemartin and d’Haultfoeuille (2020). Callaway et al. (2024) consider a treatment that has a mass at zero (constituting control group), although might also take continuously distributed non-zero doses. This permits identifying the ATET of a specific non-zero treatment dose versus no treatment, under the parallel trend assumption that the treatment group (receiving a non-zero dose) and the control group share the same trend in mean outcomes under non-treatment. However, Callaway et al. (2024) does not consider semi- or nonparametric estimation conditional on covariates. Zhang (2024) adopts a conditional parallel trends assumption in the continuous treatment context, implying that parallel trends are assumed to hold conditional on observed covariates, as previously introduced by Abadie (2005) for binary treatments. Zhang (2024) studies the canonical two-period model (with one pre- and one post-treatment period) in both the panel and repeated cross-section data, providing identification, estimation, and inference results using the double/debiased machine learning (DML) framework.
Our paper extends these results in several ways. First, we permit comparisons between two non-zero treatment doses under our stronger parallel trends assumption. Second, we allow for more than two time periods with possibly non-zero treatment doses that may vary over time, such as in a staggered adoption scenario - as e.g., considered for binary treatments in Goodman-Bacon (2021), Sant’Anna and Zhao (2020), Sun and Abraham (2021), and Borusyak et al. (2024)), which motivates conditioning on past treatment histories when assessing the ATET. Third, our approach for repeated cross-sections is capable of appropriately tackling covariate distributions that change within treatment groups over time, while many established approaches require them to be time-invariant within treatment groups, as pointed out in Hong (2013), which concerns e.g., the DID methods for binary or discrete treatments in Abadie (2005) using inverse probability weighting, Sant’Anna and Zhao (2020) using DR, or Chang (2020) using DML. A notable exception is the DML approach in Zimmert (2020), which allows for time-varying covariate distributions within treatment groups, nonetheless applying to binary rather than continuous treatments.
de Chaisemartin et al. (2024) consider treatments with strictly non-zero doses, and discuss effect identification and estimation under a parallel trend assumptions related to those considered in the present paper. The authors focus on evaluating an average ATET across the treatment margins of all subjects of changing their treatment dose, which are considered as treatment dose - e.g., what is the average effect of marginally increasing the vaccination rate on mortality among all regions that increase the vaccination rate by possibly different margins. Conversely, our study suggests estimators for assessing the ATET based on comparing two specific treatment doses rather than averaging across many doses - e.g., the effect of a vaccination rate of 40% versus 30%. As a further distinction, de Chaisemartin et al. (2024) do not consider controlling for covariates when imposing parallel trends, while our estimators permit controlling for covariates in a data-driven manner through machine learning.
The remainder of this paper is organized as follows. Section 2 discusses the DiD-based identification of the ATET with a continuous treatment for repeated cross-sectional data. Section 3 considers identification in panel data. Section 4 proposes estimators based on DML with cross-fitting. Section 5 provides asymptotic results for the proposed estimators. Section 6 presents a simulation study for the repeated cross-section and panel data cases. Section 7 concludes.
2 Identification in repeated cross-sections
This section discusses DiD-based identification in repeated cross-sectional data. Let us first consider a continuously distributed treatment, denoted by $D$, with probability mass at $D=0$. In the first case, we assume that there is a single point in time at which the treatment switches from zero to a positive treatment intensity for a subset of the population (i.e., the treated subjects), while it remains at zero for another subset which constitutes the control group. Accordingly, let $T$ denote the time period, where $T=0$ is the baseline period prior to treatment assignment, while $T>0$ is a post-treatment period after the introduction of a positive treatment dose among the treated. The treatment dose may differ across subjects, although in a first step it is assumed to remain constant across post-treatment periods. $Y_{t}$ denotes the observed outcome in period $T=t$, while $Y_{t}(d)$ is the potential outcome in period $t$ when hypothetically setting the treatment to $D=d$, see e.g., Neyman (1923) and Rubin (1974) for further discussion on the potential outcome framework. Throughout the present work, we impose the stable unit treatment value assumption (SUTVA), meaning that the potential outcomes of a subject are not affected by the treatment of others - see, for instance, the discussion in Rubin (1980) and Cox (1958). This permits defining the average treatment effect on the treated (ATET) of treatment dose $d$ in period $t$ as
$\Delta_{d,t}=E[Y_{t}(d)-Y_{t}(0)|D=d].$ (1)
Furthermore, we denote by $X$ observed covariates that may serve as control variables. We impose the following identifying assumptions, which are closely related to those in Lechner (2011), although referring to a continuous (instead of binary) treatment and multiple post-treatment periods.
Assumption 1
(Conditional parallel trends under non-treatment):
$E[Y_{t}(0)-Y_{0}(0)|D=d,X]=E[Y_{t}(0)-Y_{0}(0)|D=0,X]$
for all $d>0$ in the support of $D$ and $t>0$ in the support of $T$.
Assumption 2
(No anticipation):
$E[Y_{0}(d)-Y_{0}(0)|D=d,X]=0$
for all $d>0$ in the support of $D$.
Assumption 3
(Common support):
$\Pr(D=d,T=t|X,(D,T)\in\{(d^{\prime},t^{\prime}),(d,t)\})<1$
for all $d>0$ in the support of $D$, $t>0$ in the support of $T$, and $(d^{\prime},t^{\prime})\in\{(d,0),(0,t),(0,0)\}$.
Assumption 4
(Exogenous covariates):
$X(d)=X$
for all $d$ in the support of $D$.
Assumption 1 imposes parallel trends in the mean potential outcome under non-treatment across treated groups, receiving a non-zero treatment dose in the one hand and the control group with $D=0$ on the other hand, conditional on covariates $X$. Assumption 2 imposes that conditional on $X$, the ATET must be zero in the baseline period, which on average rules out non-zero anticipation effects among the treated in period $T=0$ regarding the treatment to be applied. Assumption 3 requires that for any treated subject in a post-treatment period $t>0$, there exist in terms of $X$ comparable treated subjects in the baseline period, non-treated subjects in the post-treatment period, and non-treated subjects in the baseline period. Lastly, Assumption 4 imposes that the covariates are not causally affected by the treatment. We note that may be a relevant restriction if they include post-treatment variables - as it is frequently the case in repeated cross-sections (see, for instance, the discussion in Caetano et al. (2022).
Under these assumptions, the conditional ATET of treatment dose $d>0$ in period $t>0$ given $X$ is identified:
$E[Y_{t}(d)-Y_{t}(0)|D=d,X]$
$=$ $E[Y_{t}(d)|D=d,X]-E[Y_{0}(0)|D=d,X]-E[Y_{t}(0)|D=d,X]+E[Y_{0}(0)|D=d,X]$
$=$ $E[Y_{t}(d)|D=d,X]-E[Y_{0}(d)|D=d,X]-\{E[Y_{t}(0)|D=d,X]-E[Y_{0}(0)|D=d,X]\}$
$=$ $E[Y_{t}(d)|D=d,X]-E[Y_{0}(d)|D=d,X]-\{E[Y_{t}(0)|D=0,X]-E[Y_{0}(0)|D=0,X]\}$
$=$ $E[Y_{t}|D=d,X]-E[Y_{0}|D=d,X]-\{E[Y_{t}|D=0,X]-E[Y_{0}|D=0,X]\}.$
The first equality in (2) follows from subtracting and adding $E[Y_{0}(0)|D=d,X]$, and the second from Assumption 2, implying that $E[Y_{0}(0)|D=d,X]=E[Y_{0}(d)|D=d,X]$. The third equality follows from Assumption 1. The fourth equality follows from the fact that $Y_{t}=Y_{t}(d)$ conditional on $D=d$, which is known as the observational rule. By Assumption 4, the effect of the treatment on the outcome $Y_{t}$ is not mediated by $X$, such that the conditional ATET in expression (2) captures the total causal effect among the treated given $X$, rather than only a partial effect not mediated by $X$. Averaging the conditional ATET over the distribution of $X$ among the treated in a post-treatment period $T=t$ yields the ATET in that period:
$\Delta_{d,t}=E[\mu_{d}(t,X)-\mu_{d}(0,X)-(\mu_{0}(t,X)-\mu_{0}(0,X))|D=d,T=t],$ (3)
where we use the short hand notation $\mu_{d}(t,x)=E[Y_{t}|D=d,X=x]$ for the conditional mean outcome given the treatment, the time period, and the covariates.
If the treatment $d$ consisted of discrete treatment doses, the ATET could alternatively be assessed based on the following doubly robust (DR) expression, which is in analogy to Zimmert (2020) for binary treatments:
$\Delta_{d,t}=E\left[\frac{I\{D=d\}\cdot I\{T=t\}\cdot[Y_{T}-\mu_{d}(0,X)-\mu_{0}(t,X)+\mu_{0}(0,X)]}{\Pi_{d,t}}\right.$
$-\left\{\frac{I\{D=d\}\cdot I\{T=0\}\cdot\rho_{d,t}(X)}{\rho_{d,0}(X)\cdot\Pi_{d,t}}\right.$ (4)
$-\left.\left.\left(\frac{I\{D=0\}\cdot I\{T=t\}\cdot\rho_{d,t}(X)}{\rho_{0,t}(X)\cdot\Pi_{d,t}}-\frac{I\{D=0\}\cdot I\{T=0\}\cdot\rho_{d,t}(X)}{\rho_{0,0}(X)\cdot\Pi_{d,t}}\right)\right\}\cdot(Y_{T}-\mu_{D}(T,X))\right],$
where $\Pi=\Pr(D=d,T=t)$ denotes the unconditional probability of receiving treatment dose $d>0$, and being observed in period $t>0$. $\rho_{d,t}(X)=\Pr(D=d,T=t|X)$ for $d\geq 0$ and $t\geq 0$ are the conditional probabilities (also known as propensity scores) of specific treatment group-period-combination $D=d,T=t$ given $X$. Lastly, $I\{\cdot\}$ denotes the indicator function, which is equal to one if its argument is satisfied and zero otherwise.
To see that the DR expression (2) corresponds to the ATET in equation (3), note that the first term on the right of equation (2), $E\left[\frac{I\{D=d\}\cdot I\{T=t\}\cdot[Y_{T}-\mu_{d}(0,X)-\mu_{0}(t,X)+\mu_{0}(0,X)]}{\Pi_{d,t}}\right]$, is by the law of itereated expectations equal to $E[\mu_{d}(t,X)-\mu_{d}(0,X)-(\mu_{0}(t,X)-\mu_{0}(0,X))|D=d,T=t]$ and thus, corresponding to the ATET:
$E\left[\frac{I\{D=d\}\cdot I\{T=t\}\cdot[Y_{T}-\mu_{d}(0,X)-\mu_{0}(t,X)+\mu_{0}(0,X)]}{\Pi_{d,t}}\right]$
$=E[Y_{T}-\mu_{d}(0,X)-(\mu_{0}(t,X)-\mu_{0}(0,X))|D=d,T=t]$
$=E[\mu_{d}(t,X)-\mu_{d}(0,X)-(\mu_{0}(t,X)-\mu_{0}(0,X))|D=d,T=t].$
Furthermore, by the law of iterated expectations and basic probability theory, we have for any $d^{\prime}$ and $t^{\prime}$ in the support of $D$ and $T$, respectively, that
$E\left[\frac{I\{D=d^{\prime}\}\cdot I\{T=t^{\prime}\}\cdot\rho_{d,t}(X)}{\rho_{d^{\prime},t^{\prime}}(X)\cdot\Pi_{d,t}}\cdot(Y_{T}-\mu_{D}(T,X))\right]$
$=E\left[\frac{\rho_{d,t}(X)}{\Pi_{d,t}}\cdot E\left[\frac{I\{D=d^{\prime}\}\cdot I\{T=t^{\prime}\}\cdot(Y_{T}-\mu_{D}(T,X))}{\rho_{d^{\prime},t^{\prime}}(X)}\Big{|}X\right]\right]$
$=E\left[\frac{\rho_{d,t}(X)}{\Pi_{d,t}}\cdot E[Y_{T}-\mu_{D}(T,X)|D=d^{\prime},T=t^{\prime},X]\right]$
$=E[\mu_{d^{\prime}}(t^{\prime},X)-\mu_{d^{\prime}}(t^{\prime},X)|D=d,T=t]=0.$
Consequently, the remaining terms on the right of equation (3) are equal to zero and expressions (2), such that equation (3) are equivalent and both yield the ATET under out identifying assumptions.
To adapt the previous results based on discrete treatments into a continuous treatment, the indicator functions for non-zero treatment doses $d$ need to be replaced by kernel weighting functions with a bandwidth approaching zero to obtain identification. Thus, we denote by $\omega(D;d,h)$ a weighting function that depends on the distance between $D$ and the non-zero reference value $d$ as well as a non-negative bandwidth parameter $h$: $\omega(D;d,h)=\frac{1}{h}\mathcal{K}\left(\frac{D-d}{h}\right)$, with $\mathcal{K}$ denoting a well-behaved kernel function. The closer $h$ is to zero, the less weight is given to larger discrepancies between $D$ and $d$. Furthermore, we note that the previous (conditional) probabilities $\Pi_{d,t}$ and $\rho_{d,t}(X)$ for $d>0$ and $t\geq 0$ correspond to (conditional) density functions if the treatment is continuous. Considering, for instance, the discussion in Fan et al. (1996), the conditional treatment density, also referred to as generalized propensity score in Imai and van Dyk (2004) and Hirano and Imbens (2005), corresponds to $\rho_{d,t}(X)=\lim_{h\rightarrow 0}E[\omega(D;d,h)\cdot I\{T=t\}|X]$. Moreover, $\Pi_{d,t}=\lim_{h\rightarrow 0}E[\omega(D;d,h)\cdot I\{T=t\}]$. In contrast, $\Pi=\Pr(D=0,T=t)$ and $\rho_{0,t}(X)$ are probabilities as adopted in equation (2), due to the fact that $D$ has a mass point at zero. Applying these modifications to DR expression (2) yields
$\Delta_{d,t}=\lim_{h\rightarrow 0}E\left[\frac{\omega(D;d,h)\cdot I\{T=t\}\cdot[Y_{T}-\mu_{d}(0,X)-\mu_{0}(t,X)+\mu_{0}(0,X)]}{\Pi_{d,t}}\right.$
$-\frac{\omega(D;d,h)\cdot I\{T=0\}\cdot\rho_{d,t}(X)\cdot\epsilon_{d,0,X}}{\rho_{d,0}(X)\cdot\Pi_{d,t}}$
$-\left.\left(\frac{I\{D=0\}\cdot I\{T=t\}\cdot\rho_{d,t}(X)\cdot\epsilon_{0,t,X}}{\rho_{0,t}(X)\cdot\Pi_{d,t}}-\frac{I\{D=0\}\cdot I\{T=0\}\cdot\rho_{d,t}(X)\cdot\epsilon_{0,0,X}}{\rho_{0,0}(X)\cdot\Pi_{d,t}}\right)\right],$ (5)
where $\epsilon_{d,t,x}=(Y_{T}-\mu_{d}(t,x))$ denotes a regression residual. It is worth mentioning that equation (3.9) in Zhang (2024) also provides an DR expression for $\Delta_{d,t}$ with continuous treatment doses $d$. However, in contrast to expression (2), the results in Zhang (2024) are not applicable to setups in which the distribution of covariates $X$ changes within treatment groups over time.
In the next step, we extend our framework to allow for time-varying treatment doses, denoting the treatment dose in period $t$ by $D_{t}$. In addition, we denote by $\mathbf{D_{t}}=\{D_{0},...,D_{t}\}$ the sequence of treatment doses up to and including period $t$. We replace Assumption 1, which imposes parallel trends on the mean potential outcome under non-treatment, with Assumption 5, which imposes parallel trends in mean potential outcomes under positive treatment doses.
Assumption 5
(Conditional parallel trends):
$E[Y_{t}(d_{t}^{\prime})-Y_{t-1}(d_{t-1}^{\prime\prime})|D_{t}=d_{t},\mathbf{D_{t-1}=d_{t-1}^{\prime\prime}},X]\\ =E[Y_{t}(d_{t}^{\prime})-Y_{t-1}(d_{t-1}^{\prime\prime})|D_{t}=d_{t}^{\prime},\mathbf{D_{t-1}=d_{t-1}^{\prime\prime}},X]$
for all $d_{t}>d_{t}^{\prime}\geq 0$ in the support of $D_{t}$, $\mathbf{d_{t-1}^{\prime\prime}}\geq\mathbf{0}$ in the support of $\mathbf{D_{t-1}}$, and $t>0$ in the support of $T$.
Assumption 5 states that for two groups that receive a higher treatment dose $d_{t}$ (treatment group) and a lower dose $d_{t}^{\prime}$ (control group), respectively, in period $t$, and both receive treatment dose $d_{t-1}^{\prime\prime}$ in period $t-1$, the trend in mean potential outcomes when receiving dose $d_{t-1}^{\prime\prime}$ in period $t-1$ and dose $d_{t}^{\prime}$ in period $t$ is equal, conditional on $X$. For the special case where $d_{t}^{\prime}=d_{t-1}^{\prime\prime}$ as e.g., considered in de Chaisemartin et al. (2024), this implies that without changing the treatment between periods $t-1$ and $t$ for any group, the treatment and control groups would experience the same trend in their mean outcomes.
Assumption 5 is stronger than Assumption 1 due to the fact that it imposes parallel trends w.r.t. non-zero treatment doses, which generally requires further assumptions, such as homogeneity in average treatment effects across groups - see e.g., Fricke (2017). To illustrate this, assume that $\mathbf{D_{t-1}}$ is a single period (instead of a sequence) $D_{t-1}$, and consider setting (1) $d_{t}^{\prime}=d_{t-1}^{\prime\prime}=0$ (such that Assumption 5 is closely related to Assumption 1) and (2) $d_{t}^{\prime}>d_{t-1}^{\prime\prime}=0$. Taking differences in the conditional means in Assumption 5 under (2) and (1) demonstrates the implied effect homogeneity across groups with $D_{t}=d$ and $D_{t}=d^{\prime}$:
$E[Y_{t}(d_{t}^{\prime})-Y_{t-1}(0)|D_{t}=d_{t},D_{t-1}=0,X]-E[Y_{t}(0)-Y_{t-1}(0)|D_{t}=d_{t},D_{t-1}=0,X]$ (6)
$=$ $E[Y_{t}(d_{t}^{\prime})-Y_{t-1}(0)|D_{t}=d_{t}^{\prime},D_{t-1}=0,X]-E[Y_{t}(0)-Y_{t-1}(0)|D_{t}=d_{t}^{\prime},D_{t-1}=0,X]$
$\leftrightarrow$ $E[Y_{t}(d^{\prime})-Y_{t}(0)|D_{t}=d_{t},D_{t-1}=0,X]=E[Y_{t}(d^{\prime})-Y_{t}(0)|D_{t}=d_{t}^{\prime},D_{t-1}=0,X].$
Furthermore, we replace the previous Assumption 2 (ruling out anticipation) by the following condition:
Assumption 6
(No anticipation):
$E[Y_{t-1}(d_{t})-Y_{t-1}(d_{t}^{\prime})|D=d_{t},\mathbf{D_{t-1}},X]=0$
for all $d_{t}\neq d_{t}^{\prime}$ in the support of $D_{t}$.
Lastly, the previous common support condition given in Assumption 3 is replaced by the following assumption, where $f(A|B)$ denotes the conditional density of $A$ given $B$:
Assumption 7
(Common support):
$f(D_{t}=d_{t},T=t|\mathbf{D_{t-1}},X,(D_{t},T)\in\{(d_{t}^{*},t^{\prime}),(d_{t},t)\})<1$
for all $d_{t}>0$ in the support of $D_{t}$, $t>0$ in the support of $T$, and $(d_{t}^{*},t^{\prime})\in\{(d_{t},t-1),(d_{t}^{\prime},t),(d_{t}^{\prime},t-1)\}$.
We note that our modified set of Assumptions 4 to 7 does not rule out that treatment doses in earlier periods affect outcomes in later periods. Thus, the treatment sequence $\mathbf{D_{t-1}}$ may influence the outcome in period $t$. In this case, we may nevertheless identify the ATET of treatment dose $d_{t}$ versus $d_{t}^{\prime}$ in period $t$ net of any direct effects of the previous history of treatment doses on the outcome in period $t$, formally defined as
$\Delta_{d_{t},d_{t}^{\prime},t}=E[Y_{t}(d_{t})-Y_{t}(d_{t}^{\prime})|D_{t}=d_{t}].$ (7)
It is worth noticing that under Assumptions 4 to 7, the conditional ATET of treatment dose $d_{t}$ versus $d^{\prime}_{t}$ in period $t$ given covariates $X$ and the previous treatment history $\mathbf{D_{t-1}}$ is analogously satisfied in expression (2), under the former Assumptions 1 to 4:
$E[Y_{t}(d_{t})-Y_{t}(d^{\prime}_{t})|D_{t}=d_{t},\mathbf{D_{t-1}},X]$
$=$ $E[Y_{t}|D_{t}=d_{t},\mathbf{D_{t-1}},X]-E[Y_{t-1}|D_{t}=d_{t},\mathbf{D_{t-1}},X]$
-- $\{E[Y_{t}|D_{t}=d^{\prime}_{t},\mathbf{D_{t-1}},X]-E[Y_{t-1}|D_{t}=d^{\prime}_{t},\mathbf{D_{t-1}},X]\}.$
Averaging the conditional ATET over the distribution of $X$ and $\mathbf{D_{t-1}}$ among the treated in any post-treatment period yields the ATET in that period:
$\Delta_{d_{t},d_{t}^{\prime},t}$ $=$ $E[\mu_{d_{t}}(t,\mathbf{D_{t-1}},X)-\mu_{d_{t}}(t-1,\mathbf{D_{t-1}},X)$
-- $(\mu_{d^{\prime}_{t}}(t,\mathbf{D_{t-1}},X)-\mu_{d^{\prime}_{t}}(t-1,\mathbf{D_{t-1}},X))|D_{t}=d_{t},T=t],$
where $\mu_{d_{t}}(t^{\prime},\mathbf{d_{t-1}},x)=E[Y_{t^{\prime}}|D_{t}=d_{t},\mathbf{D_{t-1}}=\mathbf{d_{t-1}},X=x]$ is the conditional mean outcome in period $t^{\prime}$ given the treatment received in period $t$, the previous treatment history up to period $t-1$, and the covariates. We note that for the case that $d_{t}^{\prime}=d_{t-1}^{\prime\prime}$, the ATET corresponds to a causal effect of switching from the previous treatment dose $d_{t}^{\prime}$ to a higher treatment dose $d_{t}$ among switchers. As a comparison, de Chaisemartin et al. (2024) consider average effects among switchers across different treatment margins, rather than exclusively switching to dose $d_{t}$.
As an alternative to equation (2), the ATET may be formalized by the following, equivalent DR expression:
$\Delta_{d_{t},d_{t}^{\prime},t}=$
$\lim_{h\rightarrow 0}E\left[\frac{\omega(D_{t};d_{t},h)\cdot I\{T=t\}\cdot[Y_{T}-\mu_{d_{t}}(t-1,\mathbf{D_{t-1}},X)-\mu_{d^{\prime}_{t}}(t,\mathbf{D_{t-1}},X)+\mu_{d^{\prime}_{t}}(t-1,\mathbf{D_{t-1}},X)]}{\Pi_{d_{t},t}}\right.$
$-\frac{\omega(D_{t};d_{t},h)\cdot I\{T=t-1\}\cdot\rho_{d_{t},t}(\mathbf{D_{t-1}},X)\cdot\epsilon_{d_{t},t-1,\mathbf{D_{t-1}},X}}{\rho_{d_{t},t-1}(\mathbf{D_{t-1}},X)\cdot\Pi_{d_{t},t}}$ (10)
$-\left(\frac{\omega(D_{t};d^{\prime}_{t},h)\cdot I\{T=t\}\cdot\rho_{d_{t},t}(\mathbf{D_{t-1}},X)\cdot\epsilon_{d^{\prime}_{t},t,\mathbf{D_{t-1}},X}}{\rho_{d^{\prime}_{t},t}(\mathbf{D_{t-1}},X)\cdot\Pi_{d_{t},t}}\right.$
$-\left.\left.\frac{\omega(D_{t};d^{\prime}_{t},h)\cdot I\{T=t-1\}\cdot\rho_{d_{t},t}(\mathbf{D_{t-1}},X)\cdot\epsilon_{d^{\prime}_{t},t-1,\mathbf{D_{t-1}},X}}{\rho_{d_{t}^{\prime},t-1}(\mathbf{D_{t-1}},X)\cdot\Pi_{d_{t},t}}\right)\right],$
where $\epsilon_{{d_{t}},t,\mathbf{d_{t-1}},x}=(Y_{T}-\mu_{d_{t}}(t,\mathbf{d_{t-1}},x))$ denotes a regression residual.
As a final twist to our DiD framework, we might also be interested in the ATET of a treatment dose on lagged outcomes in later periods, rather than only evaluating the the impact of a treatment dose on an outcome in the exact same period. Hence, let us denote by $s\geq 1$ the time lag between the treatment and the outcome period, which permits defining the ATET of a treatment dose in period $t-s$ on an outcome in period $t$:
$\Delta_{d_{t-s},d^{\prime}_{t-s},t}=E[Y_{t}(d_{t-s})-Y_{t}(d^{\prime}_{t-s})|D_{t-s}=d_{t-s}],$ (11)
with time lag $s\geq 1$. We note that the ATET $\Delta_{d_{t-s},d^{\prime}_{t-s},t}$ comprises both the direct effect effect of the treatment dose in period $t-s$ on the outcome in $t$, as well as the indirect (or mediated) effect of the treatment dose in period $t-s$ on the treatments in later periods, that in turn also affect the outcome in period $t-s$. This is distinct from the scenario considered in De Chaisemartin and d’Haultfoeuille (2024), where the counterfactual is defined in terms of a fixed treatment dose $d^{\prime}$ over all post treatment periods $t-s$ to $t$, such that one could, for instance, consider the ATET $E[Y_{t}(d_{t-s})-Y_{t}(d^{\prime}_{t-s},...,d^{\prime}_{t})|D_{t-s}=d_{t-s}]$, in which any indirect effects of the treatment dose in period $t-s$ are absent.
$\Delta_{d_{t-s},d^{\prime}_{t-s},t}$ is identified when replacing Assumption 5 by the following parallel trend assumption:
Assumption 8
(Conditional parallel trends with time lags):
$E[Y_{t}(d_{t-s}^{\prime})-Y_{t-s-1}(d_{t-s-1}^{\prime\prime})|D_{t-s}=d_{t-s},\mathbf{D_{t-s-1}=d_{t-s-1}^{\prime\prime}},X]\\ =E[Y_{t}(d_{t-s}^{\prime})-Y_{t-s-1}(d_{t-s-1}^{\prime\prime})|D_{t-s}=d_{t-s}^{\prime},\mathbf{D_{t-s-1}=d_{t-s-1}^{\prime\prime}},X]$
for all $d_{t-s}>d_{t-s}^{\prime}\geq 0$ in the support of $D_{t-s}$, $\mathbf{d_{t-s-1}^{\prime\prime}}\geq\mathbf{0}$ in the support of $\mathbf{D_{t-s-1}}$, $t>1$ in the support of $T$, and $1\leq s<t$.
Assumption 8 is stronger than Assumption 5, because it requires that changes in the treatment dose after period $t-s$ up to period $t$ do not evolve endogenously. This means that unobserved confounders jointly affecting the treatment doses in period $t-s$ and subsequent $s$ periods are ruled out. Otherwise, the parallel trend assumption would fail, as the outcome is also a function of the treatments after period $t-s$, such that the presence of unobserved confounders jointly affecting the treatment doses in period $t-s$ and subsequent periods would ultimately imply confounding of treatment $D_{t-s}$ and outcome $Y_{t}$. In contrast, Assumption 7 does not put restrictions on confounders of the treatment dose over time, with the caveat that the ATET may only be assessed on outcomes in the same period as the treatment dose to be evaluated, but not on outcomes in later periods.
To identify the ATET under Assumptions 4, 6, 7, and 8 using an DR expression, we just need to appropriately modify equation (2) to include time lags when defining the pre- and and post-treatment periods of interest:
$\Delta_{d_{t-s},d^{\prime}_{t-s},t}=\lim_{h\rightarrow 0}E\left[\left\{\frac{\omega(D_{t-s};d_{t-s},h)\cdot I\{T=t\}}{\Pi_{d_{t-s},t}}\right.\right.$
×$\times[Y_{T}-\mu_{d_{t-s}}(t-s-1,\mathbf{D_{t-s-1}},X)-\mu_{d^{\prime}_{t-s}}(t,\mathbf{D_{t-s-1}},X)+\mu_{d^{\prime}_{t-s}}(t-s-1,\mathbf{D_{t-s-1}},X)]$
$-\frac{\omega(D_{t-s};d_{t-s},h)\cdot I\{T=t-s-1\}\cdot\rho_{d_{t-s},t}(\mathbf{D_{t-s-1}},X)\cdot\epsilon_{d_{t-s},t-s-1,\mathbf{D_{t-s-1}},X}}{\rho_{d_{t-s},t-s-1}(\mathbf{D_{t-s-1}},X)\cdot\Pi_{d_{t-s},t}}$
$-\left(\frac{\omega(D_{t-s};d^{\prime}_{t-s},h)\cdot I\{T=t\}\cdot\rho_{d_{t-s},t}(\mathbf{D_{t-s-1}},X)\cdot\epsilon_{d^{\prime}_{t-s},t,\mathbf{D_{t-s-1}},X}}{\rho_{d^{\prime}_{t-s},t}(\mathbf{D_{t-s-1}},X)\cdot\Pi_{d_{t-s},t}}\right.$ (12)
$-\left.\left.\left.\frac{\omega(D_{t-s};d^{\prime}_{t-s},h)\cdot I\{T=t-s-1\}\cdot\rho_{d_{t-s},t}(\mathbf{D_{t-s-1}},X)\cdot\epsilon_{d^{\prime}_{t-s},t-s-1,\mathbf{D_{t-s-1}},X}}{\rho_{d_{t-s}^{\prime},t-s-1}(\mathbf{D_{t-s-1}},X)\cdot\Pi_{d_{t-s},t}}\right)\right\}\right].$
Appendix A.1 demonstrates that DR expressions (2), (2), and (2) satisfy Neyman orthogonality. Consequently, those are attractive for constructing DML estimators under certain regularity conditions outlined in Chernozhukov et al. (2018).
3 Identification in panel data
This section discusses identification in panel data under the identifying assumptions proposed in the previous section. We first reconsider the identification of $\Delta_{d_{t},d_{t}^{\prime},t}$, i.e., the ATET of treatment dose $d_{t}$ versus $d_{t}^{\prime}$ in period $t$ (net of direct effects of previous treatment doses) as defined in equation (7). As panel data permits taking outcome differences within subjects over time, the identification result in equation (2) simplifies as follows:
$\Delta_{d_{t},d_{t}^{\prime},t}$ $=$ $E[E[Y_{t}|D_{t}=d_{t},\mathbf{D_{t-1}}=\mathbf{d_{t-1}},X=x]-E[Y_{t-1}|D_{t}=d_{t},\mathbf{D_{t-1}}=\mathbf{d_{t-1}},X=x]$ (13)
-- $\{E[Y_{t}|D_{t}=d_{t}^{\prime},\mathbf{D_{t-1}}=\mathbf{d_{t-1}},X=x]$
-- $E[Y_{t-1}|D_{t}=d_{t}^{\prime},\mathbf{D_{t-1}}=\mathbf{d_{t-1}},X=x]\}|D_{t}=d_{t},T=t]$
$=$ $E[E[Y_{t}-Y_{t-1}|D_{t}=d_{t},\mathbf{D_{t-1}}=\mathbf{d_{t-1}},X=x]$
-- $\{E[Y_{t}-Y_{t-1}|D_{t}=d_{t}^{\prime},\mathbf{D_{t-1}}=\mathbf{d_{t-1}},X=x]\}|D_{t}=d_{t}].$
It is worth noticing that dropping $T=t$ from the conditioning set after the second equals sign follows from the fact that in panel data, the same subjects are observed in periods $t$ and $t-1$, such that distributions of $\mathbf{D_{t-1}}$ and $X$ are constant across periods by design.
To ease notation, we henceforth denote conditional mean differences within treatment groups across time periods by $m_{d_{t}}(t,\mathbf{D_{t-1}},X)=E[Y_{t}-Y_{t-1}|D_{t}=d_{t},\mathbf{D_{t-1}},X]$. Furthermore, we denote by $P_{d}$ and $p_{d,t}(\mathbf{D_{t-1}},X)$ (conditional) density functions of treatment $D=d$ that correspond to $P_{d_{t}}=\lim_{h\rightarrow 0}E[\omega(D;d_{t},h)]$ and $p_{d_{t}}(\mathbf{D_{t-1}},X)=\lim_{h\rightarrow 0}E[\omega(D;d_{t},h)|\mathbf{D_{t-1}},X]$, respectively. We propose an DR expression that is a straightforward modification of the one provided in equation (3.8) of Zhang (2024). The key differences are that the dose in the control group is non-zero - thus, involving kernel functions with respect to both $d_{t}^{\prime}$ and $d_{t}$, and that the conditioning set includes the treatment history $\mathbf{D_{t-1}}$ due to the multiple-period setup. Our proposed DR expression is equivalent to equation (13) for ATET identification, being given by:
$\Delta_{d_{t},d_{t}^{\prime},t}$ $=\lim_{h\rightarrow 0}E\left[\frac{\omega(D_{t};d_{t},h)\cdot[Y_{t}-Y_{t-1}-m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)]}{P_{d_{t}}}\right.$
$-\left.\frac{\omega(D_{t};d_{t}^{\prime},h)\cdot p_{d_{t}}(\mathbf{D_{t-1}},X)\cdot[Y_{t}-Y_{t-1}-m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)]}{p_{d_{t}^{\prime}}(\mathbf{D_{t-1}},X)\cdot P_{d_{t}}}\right]$ (14)
To see the equivalence, we note that by the law of iterated expectation,
$\lim_{h\rightarrow 0}E\left[\frac{\omega(D_{t};d_{t},h)\cdot[Y_{t}-Y_{t-1}-m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)]}{P_{d_{t}}}\right]$
$=E[Y_{t}-Y_{t-1}-m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)|D_{t}=d_{t}]=E[m_{d_{t}}(t,\mathbf{D_{t-1}},X)-m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)|D_{t}=d_{t}],$
which corresponds to the ATET of interest. In addition, applying basic probability theory and the law of iterated expectations, we have that
$E\left[\frac{\omega(D_{t};d_{t}^{\prime},h)\cdot p_{d_{t}}(\mathbf{D_{t-1}},X)}{p_{d_{t}^{\prime}}(\mathbf{D_{t-1}},X)\cdot P_{d_{t}}}\cdot[Y_{t}-Y_{t-1}-m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)]\right]$
$=E\left[\frac{p_{d_{t}}(\mathbf{D_{t-1}},X)}{P_{d_{t}}}\cdot E\left[\frac{\omega(D_{t};d_{t}^{\prime},h)\cdot[Y_{t}-Y_{t-1}-m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)]}{p_{d_{t}^{\prime}}(\mathbf{D_{t-1}},X)}\Big{|}X\right]\right]$
$=E\left[m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)-m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)|D_{t}=d_{t}\right]=0,$
which demonstrates that both equations (3) and (13) identify the ATET under Assumptions 4 to 7.
Lastly, we consider evaluating $\Delta_{d_{t-s},d^{\prime}_{t-s},t}$, the ATET of receiving the treatment dose $d$ versus $d^{\prime}$ in period $t-s$ on a later outcome in period $t$, as defined in equation (11). Under the satisfaction of Assumptions 4, 6, 7, and 8, the DR expression (3) is modified to the following equation:
$\Delta_{d_{t-s},d^{\prime}_{t-s},t}$ $=\lim_{h\rightarrow 0}E\left[\frac{\omega(D_{t-s};d_{t-s},h)\cdot[Y_{t}-Y_{t-s-1}-m_{d_{t-s}^{\prime}}(t,\mathbf{D_{t-s-1}},X)]}{P_{d_{t-s}}}\right.$ (15)
$-\left.\frac{\omega(D_{t-s};d_{t-s}^{\prime},h)\cdot p_{d_{t-s}}(\mathbf{D_{t-s-1}},X)\cdot[Y_{t}-Y_{t-s-1}-m_{d_{t-s}^{\prime}}(t,\mathbf{D_{t-s-1}},X)]}{p_{d_{t-s}^{\prime}}(\mathbf{D_{t-s-1}},X)\cdot P_{d_{t-s}}}\right],$
where $m_{d_{t-s}^{\prime}}(t,\mathbf{D_{t-s-1}},X)=E[Y_{t}-Y_{t-s-1}|D_{t-s}=d_{t-s},\mathbf{D_{t-s-1}},X]$. Appendix A.1 demonstrates that DR expressions (3), and (15) satisfy Neyman orthogonality.
4 Estimation
In this section, we propose DiD estimators based on DML framework with cross-fitting, as discussed in Chernozhukov et al. (2018), which are based on the sample analogs of the DR identification results in sections 2 and 3. We first consider estimating the ATET $\Delta_{d_{t},d_{t}^{\prime},t}$ as defined in equation (7) based on the sample analog of equation (2). Thus, let $\mathcal{W}=\{W_{i}|1\leq i\leq n\}$ with $W_{i}=(Y_{i,T},D_{i,t},\mathbf{D_{i,t-1}}X_{i},T_{i})$ for $i=1,\ldots,n$ denote the set of observations in an i.i.d. sample of size $n$. Moreover, let $\mathcal{W}=\{W_{i}|1\leq i\leq n\}$ with $W_{i}=(Y_{i,T},D_{i,t},\mathbf{D_{i,t-1}}X_{i},T_{i})$ for $i=1,\ldots,n$ denote the set of observations in an i.i.d. sample of size $n$. Furthermore,
$\eta=$$\{\mu_{d_{t}}(t-1,\mathbf{D_{t-1}},X),\mu_{d^{\prime}_{t}}(t,\mathbf{D_{t-1}},X),\mu_{d^{\prime}_{t}}(t-1,\mathbf{D_{t-1}},X),$
$\rho_{d_{t},t}(\mathbf{D_{t-1}},X),\rho_{d_{t},t-1}(\mathbf{D_{t-1}},X),\rho_{d^{\prime}_{t},t}(\mathbf{D_{t-1}},X),\hat{\rho}_{d^{\prime}_{t},t-1}(\mathbf{D_{t-1}},X)\}$
denotes the set of nuisance parameters, i.e., the conditional mean outcomes and treatment densities. Their respective estimates are denoted by
$\hat{\eta}=$$\{\hat{\mu}_{d_{t}}(t-1,\mathbf{D_{t-1}},X),\hat{\mu}_{d^{\prime}_{t}}(t,\mathbf{D_{t-1}},X),\hat{\mu}_{d^{\prime}_{t}}(t-1,\mathbf{D_{t-1}},X),$
$\hat{\rho}_{d_{t},t}(\mathbf{D_{t-1}},X),\hat{\rho}_{d_{t},t-1}(\mathbf{D_{t-1}},X),\hat{\rho}_{d^{\prime}_{t},t}(\mathbf{D_{t-1}},X),\hat{\rho}_{d^{\prime}_{t},t-1}(\mathbf{D_{t-1}},X)\}.$
In order to avoid overfitting, we suggest estimating $\Delta_{d_{t},d_{t}^{\prime},t}$ based on the following DML algorithm with cross-fitting, where the (kernel- and generalized propensity score-related) weights within each group are defined upon the treatment state, and the time period are normalized to sum up to one:
DML algorithm:
  1. 1.
    Split $\mathcal{W}$ in $K$ subsamples. For each subsample $k$, let $n_{k}$ denote its size, $\mathcal{W}_{k}$ the set of observations in the sample and $\mathcal{W}_{k}^{C}$ the complement set of all observations not in $\mathcal{W}_{k}$.
  2. 2.
    For each $k$, use $\mathcal{W}_{k}^{C}$ to estimate the model parameters of the nuisance parameters $\eta$ by machine learning in order to predict these nuisance parameters in $\mathcal{W}_{k}$, where the predictions are denoted by $\hat{\eta}^{k}$.
  3. 3.
    Stack the fold-specific estimates of the nuisance parameters $\hat{\eta}^{1},...,\hat{\eta}^{K}$ to generate a matrix of nuisance estimates $\hat{\eta}$ for the entire sample: $\hat{\eta}=\begin{pmatrix}\hat{\eta}^{1}\\ \vdots\\ \hat{\eta}^{K}\end{pmatrix}$
  4. 4.
    Plug the nuisance parameters into the normalized sample analog of equation (2) to obtain an estimate of the ATET using bandwidth $h$, denoted by $\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}$:
    $\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}=\sum_{i=1}^{n}\left[\biggl{(}\sum_{i=1}^{n}\omega(D_{i,t};d_{t},h)\cdot I\{T_{i}=t\}\biggr{)}^{-1}\cdot(\omega(D_{i,t};d_{t},h)\cdot I\{T_{i}=t\}\right.$
    $\cdot\left[Y_{i,T}-\hat{\mu}_{d_{t}}(t-1,\mathbf{D_{i,t-1}},X_{i})-\hat{\mu}_{d^{\prime}_{t}}(t,\mathbf{D_{i,t-1}},X_{i})+\hat{\mu}_{d^{\prime}_{t}}(t-1,\mathbf{D_{i,t-1}},X_{i})\right])$
    $-\frac{\omega(D_{i,t};d_{t},h)\cdot I\{T_{i}=t-1\}\cdot\hat{\rho}_{d_{t},t}(\mathbf{D_{i,t-1}},X_{i})\cdot\hat{\epsilon}_{i,d_{t},t-1,\mathbf{D_{i,t-1}},X_{i}}/\hat{\rho}_{d_{t},t-1}(\mathbf{D_{i,t-1}},X)}{\sum_{i=1}^{n}\omega(D_{i,t};d_{t},h)\cdot I\{T_{i}=t-1\}\cdot\hat{\rho}_{d_{t},t}(\mathbf{D_{i,t-1}},X_{i})/\hat{\rho}_{d_{t},t-1}(\mathbf{D_{i,t-1}},X)}$
    $-\left(\frac{\omega(D_{i,t};d^{\prime}_{t},h)\cdot I\{T_{i}=t\}\cdot\hat{\rho}_{d_{t},t}(\mathbf{D_{i,t-1}},X)\cdot\hat{\epsilon}_{i,d^{\prime}_{t},t,\mathbf{D_{i,t-1}},X_{i}}/\hat{\rho}_{d^{\prime}_{t},t}(\mathbf{D_{i,t-1}},X_{i})}{\sum_{i=1}^{n}\omega(D_{i,t};d^{\prime}_{t},h)\cdot I\{T_{i}=t\}\cdot\hat{\rho}_{d_{t},t}(\mathbf{D_{i,t-1}},X_{i})/\hat{\rho}_{d^{\prime}_{t},t}(\mathbf{D_{i,t-1}},X_{i})}\right.$
    $-\left.\left.\frac{\omega(D_{i,t};d^{\prime}_{t},h)\cdot I\{T_{i}=t-1\}\cdot\hat{\rho}_{d_{t},t}(\mathbf{D_{i,t-1}},X)\cdot\hat{\epsilon}_{i,d^{\prime}_{t},t-1,\mathbf{D_{i,t-1}},X_{i}}/\hat{\rho}_{d_{t}^{\prime},t-1}(\mathbf{D_{i,t-1}},X_{i})}{\sum_{i=1}^{n}\omega(D_{i,t};d^{\prime}_{t},h)\cdot I\{T_{i}=t-1\}\cdot\hat{\rho}_{d_{t},t}(\mathbf{D_{i,t-1}},X_{i})/\hat{\rho}_{d_{t}^{\prime},t-1}(\mathbf{D_{i,t-1}},X_{i})}\right)\right],$
    where $\hat{\epsilon}_{i,{d_{t}},t,\mathbf{D_{i,t-1}},X_{i}}=(Y_{i,T}-\hat{\mu}_{d_{t}}(t,\mathbf{D_{i,t-1}},X_{i}))$ denotes the estimated regression residual.
For estimation in panel data, we assume an i.i.d. sample of $n$ subjects with $W_{i}=(Y_{i,t},Y_{i,t-1},D_{i,t},\mathbf{D_{i,t-1}}X_{i})$ for $i=1,\ldots,n$. Then, the nuisance parameter vector corresponds to
$\eta=$$\{m_{d^{\prime}_{t}}(t,\mathbf{D_{t-1}},X),p_{d_{t}}(\mathbf{D_{t-1}},X),p_{d^{\prime}_{t}}(\mathbf{D_{t-1}},X)\},$
and its estimate to
$\hat{\eta}=$$\{\hat{m}_{d^{\prime}_{t}}(t,\mathbf{D_{t-1}},X),\hat{p}_{d_{t}}(\mathbf{D_{t-1}},X),\hat{p}_{d^{\prime}_{t}}(\mathbf{D_{t-1}},X)\}.$
ATET estimation proceeds analogously as outlined in the algorithm above, although using the sample analog of equation (3) in step 4 of ATET estimation:
$\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}$ $=\sum_{i=1}^{n}\left[\frac{\omega(D_{i,t};d_{t},h)\cdot[Y_{i,t}-Y_{i,t-1}-\hat{m}_{d^{\prime}_{t}}(t,t-1,\mathbf{D_{i,t-1}},X_{i})]}{\sum_{i=1}^{n}\omega(D_{i,t};d_{t},h)}\right.$
$\left.-\frac{\omega(D_{i,t};d^{\prime}_{t},h)\cdot\hat{\rho}_{d_{t}}(\mathbf{D_{i,t-1}},X)\cdot[Y_{i,t}-Y_{i,t-1}-\hat{m}_{d^{\prime}_{t}}(t,t-1,\mathbf{D_{i,t-1}},X_{i})]/\hat{\rho}_{d^{\prime}_{t}}(\mathbf{D_{i,t-1}},X_{i})}{\sum_{i=1}^{n}\omega(D_{i,t};d^{\prime}_{t},h)\cdot\hat{\rho}_{d_{t}}(\mathbf{D_{i,t-1}},X_{i})/\hat{\rho}_{d^{\prime}_{t}}(\mathbf{D_{i,t-1}},X_{i})}\right].$
Furthermore, the algorithm may be easily adapted to estimating $\Delta_{d_{t-s},d^{\prime}_{t-s},t}=E[Y_{t}(d_{t-s})-Y_{t}(d^{\prime}_{t-s})|D_{t-s}=d_{t-s}]$, the ATET of a specific treatment dose of an earlier period on a later outcome as defined in (11), by using of the normalized sample analogs of DR expressions (2) or (15) in the case of repeated cross-sections or panel data, respectively.
For estimating the variance of our estimators to conduct inference, we adapt the asymptotic variance approximation suggested in Zhang (2024) to our scenarios with non-zero treatment doses, in both the treatment and control groups. In the case of repeated cross-sections, the asymptotic variance of the ATET estimator of $\Delta_{d_{t},d_{t}^{\prime},t}$ given in (7) based on the sample analog of equation (2), denoted by $\sigma_{h}^{2}$, corresponds to
$\sigma_{h}^{2}=E\left[\left(\phi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta)-\frac{\Delta_{d_{t},d_{t}^{\prime},t}^{h}}{\Pi_{d_{t},t}}\left(\omega(D;d,h)\cdot I\{T=t\}-E[\omega(D;d,h)\cdot I\{T=t\}]\right)\right)^{2}\right].$ (16)
Defining
$\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta)$ (17)
$=\frac{\omega(D_{t};d_{t},h)\cdot I\{T=t\}\cdot[Y_{T}-\mu_{d_{t}}(t-1,\mathbf{D_{t-1}},X)-\mu_{d^{\prime}_{t}}(t,\mathbf{D_{t-1}},X)+\mu_{d^{\prime}_{t}}(t-1,\mathbf{D_{t-1}},X)]}{\Pi_{d_{t},t}}$
$-\frac{\omega(D_{t};d_{t},h)\cdot I\{T=t-1\}\cdot\rho_{d_{t},t}(\mathbf{D_{t-1}},X)\cdot\epsilon_{d_{t},t-1,\mathbf{D_{t-1}},X}}{\rho_{d_{t},t-1}(\mathbf{D_{t-1}},X)\cdot\Pi_{d_{t},t}}$ (18)
$-\left(\frac{\omega(D_{t};d^{\prime}_{t},h)\cdot I\{T=t\}\cdot\rho_{d_{t},t}(\mathbf{D_{t-1}},X)\cdot\epsilon_{d^{\prime}_{t},t,\mathbf{D_{t-1}},X}}{\rho_{d^{\prime}_{t},t}(\mathbf{D_{t-1}},X)\cdot\Pi_{d_{t},t}}\right.$
$-\left.\frac{\omega(D_{t};d^{\prime}_{t},h)\cdot I\{T=t-1\}\cdot\rho_{d_{t},t}(\mathbf{D_{t-1}},X)\cdot\epsilon_{d^{\prime}_{t},t-1,\mathbf{D_{t-1}},X}}{\rho_{d_{t}^{\prime},t-1}(\mathbf{D_{t-1}},X)\cdot\Pi_{d_{t},t}}\right),$ (19)
we have that $\Delta_{d_{t},d_{t}^{\prime},t}^{h}=E\left[\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta)\right]$ in the variance formula, which is the ATET approximation using bandwidth $h$, while $\epsilon_{{d_{t}},t,\mathbf{d_{t-1}},x}=(Y_{T}-\mu_{d_{t}}(t,\mathbf{d_{t-1}},x))$ is a regression residual. Furthermore, $\phi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta)=\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta)-\Delta_{d_{t},d_{t}^{\prime},t}^{h}$ is the Neyman-orthogonal score function under bandwidth $h$. It is worth noticing that the second term in the asymptotic variance in equation (16)
$\frac{\Delta_{d_{t},d_{t}^{\prime},t}^{h}}{\Pi_{d_{t},t}}\left(\omega(D;d,h)\cdot I\{T=t\}-E[\omega(D;d,h)\cdot I\{T=t\}]\right)$
comes for a Taylor expansion of the score w.r.t. the (low dimensional) density $P(D_{t}=d_{t},T=t)$. Therefore, it reflects estimation uncertainty related to the kernel approach required for continuous treatment doses.
A variance estimator may be constructed based on the sample analog of equation (16):
$\hat{\sigma}_{h}^{2}=\frac{1}{n}\sum_{i=1}^{n}\left[\left(\phi^{h}(W_{i},\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h},\hat{\Pi}_{k,d_{t},t},\hat{\eta})-\frac{\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}}{\hat{\Pi}_{k,d_{t},t}}\left(\omega(D;d,h)\cdot I\{T=t\}-\hat{\Pi}_{k,d_{t},t}\right)\right)^{2}\right],$ (20)
where $\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}$ is the cross-fitted ATET estimate (see step 4 of the algorithm). The term $\phi^{h}(W_{i},\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h},\hat{\Pi}_{k,d_{t},t},\hat{\eta})$ denotes the estimate of score function $\phi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta)$ for subject $i$, which is computed using the cross-fitted nuisance parameters $\hat{\eta}$ (see step 3 of the algorithm). $\hat{\Pi}_{d_{t},t}=\frac{1}{n}\sum_{i=1}^{n}\omega(D_{i};d,h)\cdot I\{T_{i}=t\}$ an estimate of the density $\Pi_{d_{t},t}$. In Zhang (2024), such types of variance estimators are shown to be consistent, provided that the kernel bandwidth $h$ satisfies $h^{-2}\varepsilon_{n}^{2}=o(1)$ for $\varepsilon_{n}=o(n^{-1/4})$.
In the case of panel data, the results in Zhang (2024) imply that the asymptotic variance of the ATET estimator of $\Delta_{d_{t},d_{t}^{\prime},t}$ takes the following form:
$\sigma_{h}^{2}=E\left[\left(\phi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)-\frac{\Delta_{d_{t},d_{t}^{\prime},t}^{h}}{P_{d_{t}}}\left(\omega(D_{t};d_{t},h)-E[\omega(D_{t};d_{t},h)]\right)\right)^{2}\right],$ (21)
where $\Delta_{d_{t},d_{t}^{\prime},t}^{h}=E[\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)]$ is the ATET approximation under bandwidth $h$ and $\phi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)=\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)-\Delta_{d_{t},d_{t}^{\prime},t}^{h}$ is the score function, with
$\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)$ $=\frac{\omega(D_{t};d_{t},h)\cdot[Y_{t}-Y_{t-1}-m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)]}{P_{d_{t}}}$ (22)
$-\frac{\omega(D_{t};d_{t}^{\prime},h)\cdot p_{d_{t}}(\mathbf{D_{t-1}},X)\cdot[Y_{t}-Y_{t-1}-m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)]}{p_{d_{t}^{\prime}}(\mathbf{D_{t-1}},X)\cdot P_{d_{t}}}.$ (23)
A cross-fitted estimator of the variance in equation (21) may be constructed in an analogous manner as outlined for the case of repeated cross-sections. Lastly, the corresponding variance expressions for the estimation of $\Delta_{d_{t-s},d^{\prime}_{t-s},t}=E[Y_{t}(d_{t-s})-Y_{t}(d^{\prime}_{t-s})|D_{t-s}=d_{t-s}]$, the ATET of a treatment dose in an earlier period on an outcome in a later period as in given (11), is obtained by a appropriately defining treatment and outcome periods according to the time lags $t-s$.
As an alternative approach to the asymptotic variance approximation, one may apply a multiplier bootstrap procedure for constructing confidence intervals. Let $\{\xi_{i}\}_{i=1}^{n}$ be an i.i.d. sequence of sub-exponential random variables independent from the data $\{W_{i}\}_{i=1}^{n}$, such that $E[\xi]=Var(\xi)=1$. In each bootstrap replication $b=1,\cdots,B$, one draws such a sequence $\{\xi_{i}\}_{i=1}^{n}$ and then constructs a bootstrap-specific ATET estimators, denoted by $\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{b,h}$, using the following expression:
$\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{b,h}=\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\cdot\psi^{h}(W_{i},\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h},\hat{\Pi}_{k,d_{t},t},\hat{\eta}).$ (24)
Denoting by $\hat{c}_{\alpha}$ the $\alpha$th quantile of the difference $\{\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{b,h}-\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}\}_{b=1}^{B}$, a $1-\alpha$ confidence interval may be constructed based on $[\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}-\hat{c}_{1-\alpha/2},\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}-\hat{c}_{\alpha/2}]$.
5 Asymptotic Theory
This section establishes the asymptotic properties of our estimators. The results largely follow Zhang (2024) and we sketch the main idea of the proof in Appendix A.2. First, we impose a set of regularity conditions on the kernel function and the boundedness and smoothness of the model parameters. These conditions are used to establish three main results. First, the bias of ATET from using a kernel approximation becomes negligible asymptotically if the kernel bandwidth is undersmoothed. Second, the ATET estimators are asymptotically normal. Third, the variance estimators based on the asymptotic variance formulas proposed in Section 4 are consistent. We focus on the case without time lags, and the results for the case with time lags follow immediately by appropriately defining treatment and outcome periods according to the time lag $t-s$. Recalling that the weighting function is defined as $\omega(D;d,h)=\frac{1}{h}\mathcal{K}\left(\frac{D-d}{h}\right)$ with $h$ being the bandwidth, we consider bandwidth-dependent versions of the generalized propensitiy scores, denoted by $\rho_{d,t}^{h}(\mathbf{D_{t-1}},X)=E[\omega(D;d,h)\cdot I\{T=t\}|\mathbf{D_{t-1}},X]$ and $p_{d}^{h}(\mathbf{D_{t-1}},X)=E[\omega(D;d,h)|\mathbf{D_{t-1}},X]$, in some the subsequent assumptions. Furthermore, we denote by $\hat{\rho}_{d,t}^{h}(\mathbf{D_{t-1}},X)$ and $\hat{p}_{d}^{h}(\mathbf{D_{t-1}},X)$ estimates of $\rho_{d,t}^{h}(\mathbf{D_{t-1}},X)$ and $p_{d}^{h}(\mathbf{D_{t-1}},X)$.
Assumption 9
(Kernel):
[…]
Assumption 10
(Bounds and smoothness, repeated cross-sections):
(a) For some constants $10<c<$ 1 and $0<C<\infty$, $\Pi_{d_{t},t}>c$, $|\rho_{d_{t},t-1}^{h}(\mathbf{D_{t-1}},X)|>c$, $|\rho_{d_{t}^{\prime},t}^{h}(\mathbf{D_{t-1}},X)|>c$, $|\rho_{d_{t}^{\prime},t-1}^{h}(\mathbf{D_{t-1}},X)|>c$, $|Y_{T}|<C$, $|\rho_{d_{t},t}^{h}(\mathbf{D_{t-1}},X)|<C$, $|\mu_{d_{t}}(t-1,\mathbf{D_{t-1}},X)|<C$, $|\mu_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)|<C$, and $|\mu_{d_{t}^{\prime}}(t-1,\mathbf{D_{t-1}},X)|<C$ almost surely; (b) $\Pi_{d_{t},t}\in C^{2}(\text{supp}(D_{t}))$ and $|\Pi_{d_{t},t}^{(2)}|<\infty$; (c) $\rho_{d_{t},t-1}(\mathbf{D_{t-1}},X)\in C^{2}(\text{supp}(D_{t}))$ and $\sup_{x\in\text{supp}(X)}|\rho_{d_{t},t-1}^{(2)}(\mathbf{D_{t-1}},X)|<\infty$; (d) joint density $f_{Y_{T},D_{t}}(y,d_{t})\in C^{2}(\text{supp}(Y_{T}))$ and $\sup_{d_{t}\in\text{supp}(D_{t})}|f_{Y_{T},D_{t}}^{(2)}(y,d_{t})|<\infty$.
Assumption 11
(Rates, repeated cross-sections):
(a) The kernel bandwidth $h$ is a deterministic sequence that depends on $n$ and satisfies $nh\to\infty$ and $\sqrt{nh^{5}}=o(1)$; (b) there exists a sequence $\varepsilon_{n}\to 0$, such that $h^{-1}\varepsilon_{n}^{2}=o(1)$; (c) with probability tending to $1$, $\|\hat{\rho}_{d_{t},t}^{h}(\mathbf{D_{t-1}},X)-\rho_{d_{t},t}^{h}(\mathbf{D_{t-1}},X)\|_{P,2}\leq h^{-1/2}\varepsilon_{n}$, $\|\hat{\rho}_{d_{t},t-1}^{h}(\mathbf{D_{t-1}},X)-\rho_{d_{t},t-1}^{h}(\mathbf{D_{t-1}},X)\|_{P,2}\leq\varepsilon_{n}$, $\|\hat{\rho}_{d_{t}^{\prime},t-1}^{h}(\mathbf{D_{t-1}},X)-\rho_{d_{t}^{\prime},t-1}^{h}(\mathbf{D_{t-1}},X)\|_{P,2}\leq\varepsilon_{n}$, $\|\hat{\mu}_{d_{t}}(t-1,\mathbf{D_{t-1}},X)-\mu_{d_{t}}(t-1,\mathbf{D_{t-1}},X)\|_{P,2}\leq\varepsilon_{n}$, $\|\hat{\mu}_{d_{t}^{\prime}}(t-1,\mathbf{D_{t-1}},X)-\mu_{d_{t}^{\prime}}(t-1,\mathbf{D_{t-1}},X)\|_{P,2}\leq\varepsilon_{n}$, $\|\hat{\mu}_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)-\mu_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)\|_{P,2}\leq\varepsilon_{n}$; (d) with probability tending to $1$, $\sup_{\text{supp}(\mathbf{D_{t-1}},X)}\hat{\rho}_{d_{t},t-1}^{h}(\mathbf{D_{t-1}},X)>c$, $\sup_{\text{supp}(\mathbf{D_{t-1}},X)}\hat{\rho}_{d_{t}^{\prime},t}^{h}(\mathbf{D_{t-1}},X)>c$, $\sup_{\text{supp}(\mathbf{D_{t-1}},X)}\hat{\rho}_{d_{t}^{\prime},t-1}^{h}(\mathbf{D_{t-1}},X)>c$, $\|\hat{\rho}_{d_{t},t}^{h}(\mathbf{D_{t-1}},X)\|_{P,\infty}<C$, and $\|\hat{\mu}_{d_{t}}(t-1,\mathbf{D_{t-1}},X)\|_{P,\infty}<C$, $\|\hat{\mu}_{d_{t}^{\prime}}(t-1,\mathbf{D_{t-1}},X)\|_{P,\infty}<C$, $\|\hat{\mu}_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)\|_{P,\infty}<C$.
Assumption 12
(Bounds and smoothness, panel):
(a) For some constants $10<c<$ 1 and $0<C<\infty$, $P_{d_{t}}>c$, $|Y_{t-1}|<C$, $|Y_{t}|<C$, $p_{d_{t}^{\prime}}^{h}(\mathbf{D_{t-1}},X)>c$, $|p_{d_{t}}^{h}(\mathbf{D_{t-1}},X)|<C$, and $|m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)|<C$ almost surely for all $h>0$; (b) $P_{d_{t}}\in C^{2}(\text{supp}(D_{t}))$ and $|P_{d_{t}}^{(2)}|<\infty$; (c) $p_{d_{t}}(\mathbf{D_{t-1}},X)\in C^{2}(\text{supp}(D_{t}))$ and $\sup_{\text{supp}(\mathbf{D_{t-1}},X)}|p_{d_{t}}^{(2)}(\mathbf{D_{t-1}},X)|<\infty$; (d) for $\Delta Y=Y_{t}-Y_{t-1}$, the joint density $f_{\Delta Y,D_{t}}(s,d_{t})\in C^{2}(\text{supp}(\Delta Y))$ and $\sup_{d\in\text{supp}(D_{t})}|f_{\Delta Y,D_{t}}^{(2)}(s,d_{t})|<\infty$.
Assumption 13
(Rates, panel):
(a) The kernel bandwidth $h$ is a deterministic sequence that depends on $n$ and satisfies $nh\to\infty$ and $\sqrt{nh^{5}}=o(1)$; (b) there exists a sequence $\varepsilon_{n}\to 0$ such that $h^{-1}\varepsilon_{n}^{2}=o(1)$; (c) with probability tending to $1$, $\|\hat{p}_{d_{t}}^{h}(\mathbf{D_{t-1}},X)-p_{d_{t}}^{h}(\mathbf{D_{t-1}},X)\|_{P,2}\leq h^{-1/2}\varepsilon_{n}$, $\|\hat{p}_{d_{t}^{\prime}}^{h}(\mathbf{D_{t-1}},X)-p_{d_{t}^{\prime}}^{h}(\mathbf{D_{t-1}},X)\|_{P,2}\leq\varepsilon_{n}$, $\|\hat{m}_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)-m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)\|_{P,2}\leq\varepsilon_{n}$; (d) with probability tending to $1$, $\sup_{\text{supp}(\mathbf{D_{t-1}},X)}\hat{p}_{d_{t}^{\prime}}^{h}(\mathbf{D_{t-1}},X)>c$, $\|\hat{p}_{d_{t}}^{h}(\mathbf{D_{t-1}},X)\|_{P,\infty}<C$, $\|\hat{m}_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)\|_{P,\infty}<C$.
Assumption 9 concerning the kernel function is standard and many commonly used smoothing functions satisfy these conditions, like the Gaussian or Epanechnikov kernel. Furthermore, we impose two sets of regularity conditions in both repeated cross-sections and panel settings. Firstly, we require that bandwidth-dependent nuisance parameters in the population are bounded and satisfy certain smoothness conditions, see Assumptions 10 and 12. Secondly, we impose high-level assumptions on the quality of the estimators of the nuisance parameters, permitting those estimators to converge at a slower rate compared to a setting without Neyman orthogonality, see Assumptions 11 and 13. A first result obtained from these conditions is that the biases in repeated cross-sections and panel settings are of order $O(h^{2})$, such that they vanish asymptotically when using an undersmoothed kernel bandwidth for ATET estimation.
Lemma 1
Suppose Assumptions 9, 10, 11 hold for the repeated cross-sections case, and Assumptions 9, 12, 13 hold for the panel case. Then, the estimation bias is given by $B(h)=\Delta_{d_{t},d_{t}^{\prime},t}-\Delta_{d_{t},d_{t}^{\prime},t}^{h}$ for $d_{t},d_{t}^{\prime}\in\text{supp}(D_{t})$.
The next two theorems are the main results of this section. The first theorem states that our proposed ATET estimators are asymptotically normal and converge at $\sqrt{nh}$ rate to the true ATET, while the second theorem states that our variance estimators proposed in the previous section are consistent.
Theorem 1
Suppose Assumptions 4, 5, 6, 7 hold. Moreover, suppose Assumptions 9, 10, 11 hold for the repeated cross-sections case, and Assumptions 9, 12, 13 hold for the panel case. Then the ATET estimators defined in Section 4 satisfy
$\frac{\hat{\Delta}_{d_{t},d_{t}^{\prime},t}-\Delta_{d_{t},d_{t}^{\prime},t}}{\sigma_{h}/\sqrt{n}}\quad\rightarrow^{d}\quad N(0,1)$ (25)
where $\sigma_{h}$ is defined as in (16) and (21) for the repeated cross-sections case and panel case, respectively.
Theorem 2
Assume Theorem 1 holds and assume that $h^{-2}\varepsilon_{n}^{2}=o(1)$. Then the variance estimators defined in Section 4 are consistent for the asymptotic variance $\sigma_{h}$ defined in (16) and (21) for the repeated cross-sections case and panel case, respectively.
A sketch of the proofs is provided in Appendix A.2.
6 Simulation study
This section provides a simulation study to investigate the finite sample behavior of our DiD estimators for repeated cross-sections and panel data. Starting with the case of repeated cross-sections, we consider the following data generating process (DGP):
$X_{j}$ $=$ $0.5T+Q_{j}\textrm{ for }j\in\{1,...,p\},\quad X=(X_{1},...,X_{p}),$
$D$ $=$ $X\beta+0.5\cdot U+V,$
$Y_{T}$ $=$ $X\beta+(1+D^{2})\cdot T+U+W,$
$Q_{j},U,V,W$ $\sim$ $Unif(0,2)\textrm{ independent of each other},$
$T$ $\sim$ $binomial(0.5).$
The vector $X$ consists of time varying covariates $X_{j}$ with $j\in\{1,...,p\}$, where $p$ is the dimension of $X$. Each covariate $X_{j}$ depends on the time index $T$, which is either 0 for the pre-treatment and 1 for the post-treatment period (each with a probability of 50%), as well as the random error $Q_{j}$, which is independently uniformly distributed with its support ranging from 0 to 2.
The continuous treatment $D$ is a function of covariates $X$ if the coefficient vector $\beta$ is different from zero and two uniformly distributed error terms, namely the fixed effect $U$ and a random component $V$. Both the covariates $X$ and the fixed effect $U$ also affect the outcome $Y_{T}$ and are, therefore, confounders of the treatment-outcome relation that is tackled by our covariate-adjusted DiD approach. In addition, the outcome is affected by the period indicator $T$, which reflects a time trend, and a random uniformly distributed error term $W$. Lastly, the treatment has a nonlinear effect on the post-treatment outcome, which is modeled by the interaction of $D^{2}$ and $T$.
In our simulations, we set the number of covariates to $p=100$ and the $j$th element in the coefficient vector $\beta$ to $0.4/i^{2}$ for $j\in\{1,...,p\}$, which implies a quadratic decay of covariate importance in terms of confounding. We consider two sample sizes of $n=2000,4000$. DiD estimation is based on the sample analog of equation (2). In our simulation design, there are only two periods $t$ such that $T=t$ and $T=t-1$ in equation (2) correspond to $T=1$ and $T=0$, respectively. Furthermore, there is no history of previous treatments in $T=0$ such that $\mathbf{D_{t-1}}=\emptyset$ in our simulations. However, from an econometric perspective, there is no difference between controlling for the treatment history $\mathbf{D_{t-1}}$ or covariates $X$, which are both time-varying sets of confounders. For this reason, not considering $\mathbf{D_{t-1}}$ in our simulations comes without loss of generality as we could easily relabel some elements in $X$ to represent $\mathbf{D_{t-1}}$.
For estimation, we apply the didcontDML function of the causalweight package by Bodory and Huber (2018) for the statistical software R (R Core Team (2020)). We run this DML-based DiD estimator for repeated cross-sections with 3-fold cross-fitting and estimate the nuisance parameters by cross-validated lasso regression. We adopt the second order Epanechnikov kernel as the weighting function $\omega$ of the continuous treatment, where the bandwidth corresponds to $2.34n^{-0.25}$. The latter choice corresponds to a Silverman (1986)-type rule of thumb for Epanechnikov kernels, as for instance also applied in Huber et al. (2020). In addition to estimation based on the rule of thumb, we consider an undersmoothed version, in which the kernel bandwidth is divided by 2. Furthermore, for estimating the generalized propensity scores $\rho_{d,t}(X)$ by lasso, we consider both linear and log-linear models that assume normally distributed unobservables $V$ (which is incorrect as $V$ is uniformly distributed). The ATET estimator also imposes a trimming rule that safeguards against disproportionately influential observations in the inverse propensity weighting (IPW) by dropping observations receiving a weight that is larger than 10% in the IPW-based computation of any conditional mean outcome given the treatment state and time period. Such a trimming rule for ensuring common support in the sample is considered in Huber et al. (2013). For obtaining the standard error of the ATET estimate, we consider the estimated asymptotic variance approximation provided in equation (20) of Section 4.
The upper panel of Table 1 provides the simulation results for repeated cross-sections, when estimating the ATET $\Delta_{3,2,1}=E[Y_{1}(3)-Y_{0}(2)|D_{1}=3]=3^{2}-2^{2}=5$, based on linear lasso regression without undersmoothing (‘lasso’), loglinear lasso regression without undersmoothing (‘lnorm’), linear lasso regression with undersmoothing (‘under’), and loglinear lasso regression with undersmoothing (‘ln under’). The columns provide the estimation method (‘method’), the bias of the estimator (‘bias’), its standard deviation (‘std’), its root mean squared error (‘rmse’), the average standard error (‘avse’), and the coverage rate of the 95% confidence intervals (‘cover’), respectively, for either sample size.
For $n=2000$, we observe that the ATET estimators without undersmoothing are non-negligibly biased, while those with undersmoothing are less biased at the price of having a somewhat higher variance. Overall, the undersmoothed versions provide a superior performance in terms of the root mean squared error (RMSE), which measures the overall estimation error as a combination of both bias and variance. Furthermore, the average standard error is generally close to the actual standard deviation, suggesting that asymptotic variance estimation as outlined in Section 4 works exceedingly well in our simulations. For the undersmoothed ATET estimators, this implies a reasonable coverage rate that is approximately the nominal level of 95%. Regarding the the estimators without undersmoothing, the coverage is notably low due to larger biases. The results are virtually not affected by the assumption of normal or lognormal models for the estimation of the generalized propensity scores. When increasing the sample size to $n=8000$, the biases of the estimators without undersmoothing are reduced, although still considerably different from zero. In contrast, the biases of the estimators with undersmoothing are negligibly small, being substantially better in terms of RMSE compared to the estimation without undersmoothing. In terms of coverage, the estimators without undersmoothing attain roughly 83%, while the undersmoothed versions attain 97%, which is nearly the nominal level.
Table 1: Simulation results
repeated cross-sections
$n=2000$ $n=8000$
method bias std rmse avse cover bias std rmse avse cover
lasso -0.376 0.249 0.451 0.251 0.682 -0.173 0.182 0.252 0.183 0.843
lnorm -0.379 0.251 0.454 0.251 0.668 -0.176 0.182 0.253 0.184 0.840
under -0.090 0.283 0.297 0.287 0.949 -0.027 0.209 0.211 0.211 0.966
ln under -0.095 0.283 0.298 0.287 0.952 -0.031 0.210 0.212 0.211 0.967
panel data
$n=2000$ $n=8000$
method bias std rmse avse cover bias std rmse avse cover
lasso -0.541 0.123 0.555 0.127 0.009 -0.220 0.056 0.227 0.055 0.023
lnorm -0.545 0.122 0.558 0.127 0.009 -0.224 0.056 0.231 0.055 0.021
under -0.067 0.103 0.123 0.111 0.921 0.039 0.058 0.070 0.061 0.910
ln under -0.071 0.103 0.125 0.111 0.916 0.034 0.058 0.067 0.060 0.922
Notes: columns ‘method’, ‘bias’, ‘std’, ‘rmse’, ‘avse’, and ‘cover’ provide estimation method, the bias of the estimator, its standard deviation, its root mean squared error, the average of the standard error, and the coverage rate of the 95% confidence intervals, respectively. The ‘lasso’ provides the results for lasso-based estimation when assuming a linear treatment model with normally distributed errors. The‘lnorm’ assumes instead a loglinear treatment model with normal errors. The ‘under’ relies on a linear treatment model with undersmoothing, using half of the bandwidth suggested by the rule of thumb for smoothing the continuous treatment. The ‘ln under’ provides the results for a loglinear treatment model and undersmoothing. The true ATET is 5.
Subsequently, we investigate the DiD estimator for panel data, considering a slightly modified DGP in which $X_{j}$ only affects the outcome in the second period, but not in the first period:
$X$ $=$ $(X_{1},...,X_{p}),$
$D$ $=$ $X\beta+0.5\cdot U+V,$
$Y_{T}$ $=$ $(1+D^{2}+X\beta)\cdot T+U+W,$
$X_{j},U,V,W$ $\sim$ $Unif(0,2)\textrm{ independent of each other and for }j\in\{1,...,p\},$
$T$$\in$ $\{0,1\}.$
We apply the didcontDMLpanel function of the causalweight package by Bodory and Huber (2018) and consider the same choices in terms of kernel functions, bandwidth selection, undersmoothing, generalized propensity score modeling, and trimming as for the repeated cross-sections case.
The lower panel of Table 1 provides the simulation results for the panel case, when again estimating the ATET $\Delta_{3,2,1}=5$. The biases of the ATET estimators without undersmoothing are once more nonnegligible under either sample size, while undersmoothing drives the bias to approach zero fast as the sample size increases, which entails a substantially superior performance in terms of RMSE. The average standard errors based on asymptotic variance approximation as outlined in Appendix 21 closely match the actual standard deviations of the respective estimators. Finally, we observe a slight under-coverage by the confidence intervals of the undersmoothed estimators. However, this is not substantially below the nominal level of 95%. In contrast, coverage is poor for the ATET estimators without undersmoothing, due to their large biases. For this reason, the simulation results for the panel and repeated cross-section cases are qualitatively fairly similar.
7 Conclusion
In this paper, we introduced the kernel-based treatment effect estimators for continuously distributed and possibly time-varying treatment doses in difference-in-differences (DiD) models. It adopts double machine learning to control for covariates and past treatment histories in a data-adaptive manner. We proposed methods for both repeated cross-sections, allowing covariate distributions within treatment groups to change over time (in contrast to most of the DiD literature), and panel data, along with asymptotic formulae for variance estimation. We also presented a simulation study, which indicates a decent finite sample performance of undersmoothed versions of our estimators in setups with several thousand observations. Our estimators are available in the causalweight for the statistical software R.
Appendix A Appendix
A.1 Neyman orthogonality of the DR expressions
We first show the Neyman orthogonality of the DR expressions (2), (2), and (2) for repeated cross-sections. We consider the ATET $\Delta_{d,t}=E[Y_{t}(d)-Y_{t}(0)|D=d]$ provided in equation (1), which is identified under Assumptions 1 to 4. In the subsequent discussion, all bounds hold uniformly over all probability laws $P\in\mathcal{P}$, where $\mathcal{P}$ is the set of all possible probability laws - we omit $P$ for brevity. We define the nuisance parameters
$\eta=(\mu_{d^{\prime}}(t^{\prime},X),\rho_{d,t}(X),\rho_{d^{\prime},t^{\prime}}(X))$
and denote the true values of the nuisance parameters in the population by
$\eta^{0}=(\mu^{0}_{d^{\prime}}(t^{\prime},X),\rho^{0}_{d,t}(X),\rho^{0}_{d^{\prime},t^{\prime}}(X)).$
We also define $\Psi_{d^{\prime},t^{\prime}}=E[\mu_{d^{\prime}}(t^{\prime},X)|D=d,T=t]$ to be the average of the conditional mean $\mu_{d^{\prime}}(t^{\prime},X)$ in treatment group $D=d$ and time period $T=t$. We note that the ATET provided in equation (3) may be expressed as a function of $\Psi_{d^{\prime},t^{\prime}}$:
$\Delta_{d,t}=E[\mu_{d}(t,X)-\mu_{d}(0,X)-(\mu_{0}(t,X)-\mu_{0}(0,X))|D=d,T=t]$
$=E[Y_{T}|D=d,T=t]-\Psi_{d,0}-\Psi_{0,t}+\Psi_{0,0},$ (A.1)
where the second equality follows from the law of iterated expectations. Therefore, considering Neyman-orthogonal functions for $\Psi_{d^{\prime},t^{\prime}}$ yields a Neyman-orthogonal expression of the ATET, which is an additive function of $\Psi_{d^{\prime},t^{\prime}}$ for $d^{\prime}$in $\{d,0\}$ and $t^{\prime}$in $\{t,0\}$.
We denote by $\phi_{d^{\prime},t^{\prime}}$ the Neyman-orthogonal score function for $\Psi_{d^{\prime},t^{\prime}}$, which corresponds to the following expression for a continuous treatment $D$ when considering values $d$ and $d^{\prime}$:
$\phi_{d^{\prime},t^{\prime}}$ $=\frac{\omega(D;d,h)\cdot I\{T=t\}\cdot\mu_{d^{\prime}}(t^{\prime},X)}{\Pi_{d,t}}$
$+\frac{\omega(D;d^{\prime},h)\cdot I\{T=t^{\prime}\}\cdot\rho_{d,t}(X)\cdot(Y_{T}-\mu_{d^{\prime}}(t^{\prime},X))}{\rho_{d^{\prime},t^{\prime}}(X)\cdot\Pi_{d,t}}-\Psi_{d^{\prime},t^{\prime}}.$ (A.2)
We note that if $d$ and/or $d^{\prime}$ are mass points (e.g., $d^{\prime}=0$), then the kernel functions $\omega(D;d,h)$ and/or $\omega(D;d^{\prime},h)$ should be replaced by indicator functions $I\{D=d\}$ and/or $I\{D=d^{\prime}\}$. For the sake of simplicity, we omit this case in our discussion and focus on the continuous case. We also note that DR expression (2) is obtained by identifying $\Psi_{d^{\prime},m^{\prime}}$ as the solution to the moment condition $E[\phi_{d^{\prime},m^{\prime}}]=0$, and plugging it into equation (A.1).
The Gateaux derivative of $E[\phi_{d^{\prime},t^{\prime}}]$ in the direction of $[\eta^{0}-\eta]$ is given by:
$\partial E[\phi_{d^{\prime},t^{\prime}}][\eta-\eta^{0}]=\lim_{h\rightarrow 0}E\left[\frac{\omega(D;d,h)\cdot I\{T=t\}\cdot[\mu_{d^{\prime}}(t^{\prime},X)-\mu^{0}_{d^{\prime}}(t^{\prime},X)]}{\Pi_{d,t}}\right]$
$-\lim_{h\rightarrow 0}E\left[\frac{\omega(D;d^{\prime},h)\cdot I\{T=t^{\prime}\}\cdot\rho^{0}_{d,t}(X)\cdot[\mu_{d^{\prime}}(t^{\prime},X)-\mu^{0}_{d^{\prime}}(t^{\prime},X)]}{\rho^{0}_{d^{\prime},t^{\prime}}(X)\cdot\Pi_{d,t}}\right]$
$+\lim_{h\rightarrow 0}E\left[\frac{\omega(D;d^{\prime},h)\cdot I\{T=t^{\prime}\}\cdot[Y_{T}-\mu^{0}_{d^{\prime}}(t^{\prime},X)]}{\rho^{0}_{d^{\prime},t^{\prime}}(X)}\cdot\frac{[\rho_{d,t}(X)-\rho^{0}_{d,t}(X)]}{\Pi_{d,t}}\right]$ (A.3)
$-\lim_{h\rightarrow 0}E\left[\frac{\omega(D;d^{\prime},h)\cdot I\{T=t^{\prime}\}\cdot\rho^{0}_{d,t}(X)\cdot[Y_{T}-\mu^{0}_{d^{\prime}}(t^{\prime},X)]}{\rho^{0}_{d^{\prime},t^{\prime}}(X)\cdot\Pi_{d,t}}\cdot\frac{\rho_{d^{\prime},t^{\prime}}(X)-\rho^{0}_{d^{\prime},t^{\prime}}(X)}{\rho^{0}_{d^{\prime},t^{\prime}}(X)}\right]=0.$
Applying the law of iterated expectations to the term in the second line of equation (A.1) yields:
$-\lim_{h\rightarrow 0}E\left[E\left[\frac{\omega(D;d^{\prime},h)\cdot I\{T=t^{\prime}\}\cdot\rho^{0}_{d,t}(X)\cdot[\mu_{d^{\prime}}(t^{\prime},X)-\mu^{0}_{d^{\prime}}(t^{\prime},X)]}{\rho^{0}_{d^{\prime},t^{\prime}}(X)\cdot\Pi_{d,t}}\ \bigg{|}X\right]\right]$
$=-\lim_{h\rightarrow 0}E\left[\frac{E[\omega(D;d^{\prime},h)\cdot I\{T=t^{\prime}\}|X]\cdot\rho^{0}_{d,t}(X)\cdot[\mu_{d^{\prime}}(t^{\prime},X)-\mu^{0}_{d^{\prime}}(t^{\prime},X)]}{\rho^{0}_{d^{\prime},t^{\prime}}(X)\cdot\Pi_{d,t}}\right]$
$=-E\left[\frac{\rho^{0}_{d^{\prime},t^{\prime}}(X)\cdot\rho^{0}_{d,t}(X)\cdot[\mu_{d^{\prime}}(t^{\prime},X)-\mu^{0}_{d^{\prime}}(t^{\prime},X)]}{\rho^{0}_{d^{\prime},t^{\prime}}(X)\cdot\Pi_{d,t}}\right]$
$=-E\left[\frac{\rho^{0}_{d,t}(X)\cdot[\mu_{d^{\prime}}(t^{\prime},X)-\mu^{0}_{d^{\prime}}(t^{\prime},X)]}{\Pi_{d,t}}\right].$
Applying the law of iterated expectations to the term on the left hand side of the first line yields:
$\lim_{h\rightarrow 0}E\left[\frac{E[\omega(D;d,h)\cdot I\{T=t\}|X]\cdot[\mu_{d^{\prime}}(t^{\prime},X)-\mu^{0}_{d^{\prime}}(t^{\prime},X)]}{\Pi_{d,t}}\right]$
$=$ $E\left[\frac{\rho^{0}_{d,t}(X)\cdot[\mu_{d^{\prime}}(t^{\prime},X)-\mu^{0}_{d^{\prime}}(t^{\prime},X)]}{\Pi_{d,t}}\right].$
Consequently, the terms in the first and in second line of equation (A.1) cancel out. Considering the third line, we have that:
$\lim_{h\rightarrow 0}E\left[E\left[\frac{\omega(D;d^{\prime},h)\cdot I\{T=t^{\prime}\}\cdot[Y_{T}-\mu^{0}_{d^{\prime}}(t^{\prime},X)]}{\rho^{0}_{d^{\prime},t^{\prime}}(X)}\cdot\frac{[\rho_{d,t}(X)-\rho^{0}_{d,t}(X)]}{\Pi_{d,t}}\bigg{|}X\right]\right]$
$=$ $\lim_{h\rightarrow 0}E\left[\frac{E[\omega(D;d^{\prime},h)\cdot I\{T=t^{\prime}\}\cdot[Y_{T}-\mu^{0}_{d^{\prime}}(t^{\prime},X)]|X]}{\rho^{0}_{d^{\prime},t^{\prime}}(X)}\cdot\frac{[\rho_{d,t}(X)-\rho^{0}_{d,t}(X)]}{\Pi_{d,t}}\right].$
Concerning the first factor, we have by basic probability theory that:
$\frac{\lim_{h\rightarrow 0}E[\omega(D;d^{\prime},h)\cdot I\{T=t^{\prime}\}\cdot Y_{T}|X]}{\rho^{0}_{d^{\prime},t^{\prime}}(X)}=E[Y_{T}|D=d^{\prime},T=t^{\prime},X]=\mu^{0}_{d^{\prime}}(t^{\prime},X).$
Furthermore,
$\frac{\lim_{h\rightarrow 0}E[\omega(D;d^{\prime},h)\cdot I\{T=t^{\prime}\}\cdot\mu^{0}_{d^{\prime}}(t^{\prime},X)|X]}{\rho^{0}_{d^{\prime},t^{\prime}}(X)}$
$=$ $\frac{\lim_{h\rightarrow 0}E[\omega(D;d^{\prime},h)\cdot I\{T=t^{\prime}\}|X]\cdot\mu^{0}_{d^{\prime}}(t^{\prime},X)}{\rho^{0}_{d^{\prime},t^{\prime}}(X)}=\frac{\rho^{0}_{d^{\prime},t^{\prime}}(X)\cdot\mu^{0}_{d^{\prime}}(t^{\prime},X)}{\rho^{0}_{d^{\prime},t^{\prime}}(X)}=\mu^{0}_{d^{\prime}}(t^{\prime},X).$
This implies that the first factor in the third line of equation (A.1) is zero, such that the entire expression in the third line is also zero. In analogous manner, we show for the fourth line that:
$-\lim_{h\rightarrow 0}E\left[E\left[\frac{\omega(D;d^{\prime},h)\cdot I\{T=t^{\prime}\}\cdot\rho^{0}_{d,t}(X)\cdot[Y_{T}-\mu^{0}_{d^{\prime}}(t^{\prime},X)]}{\rho^{0}_{d^{\prime},t^{\prime}}(X)\cdot\Pi_{d,t}}\cdot\frac{\rho_{d^{\prime},t^{\prime}}(X)-\rho^{0}_{d^{\prime},t^{\prime}}(X)}{\rho^{0}_{d^{\prime},t^{\prime}}(X)}\bigg{|}X\right]\right]$
$=-E\left[\frac{\rho^{0}_{d,t}(X)\cdot[\mu^{0}_{d^{\prime}}(t^{\prime},X)-\mu^{0}_{d^{\prime}}(t^{\prime},X)]}{\Pi_{d,t}}\cdot\frac{\rho_{d^{\prime},t^{\prime}}(X)-\rho^{0}_{d^{\prime},t^{\prime}}(X)}{\rho^{0}_{d^{\prime},t^{\prime}}(X)}\right].$
As the first factor is zero due to the fact that $[\mu^{0}_{d^{\prime}}(t^{\prime},X)-\mu^{0}_{d^{\prime}}(t^{\prime},X)]=0$, the expression in the fourth line of equation (A.1) is also zero. Therefore, the Gateaux derivative of $E[\phi_{d^{\prime},t^{\prime}}]$ in the direction of $[\eta^{0}-\eta]$ is zero. Moreover, we trivially have that the Gateaux derivative of $E[Y_{T}|D=d,T=t]$ in the direction of $[\eta^{0}-\eta]$ is zero. We note that the DR expression (2) is equivalent to expression (A.1) when using the Neyman-orthogonal score functions (A.1) for computing $\Psi_{d,0}$, $\Psi_{0,t}$ $\Psi_{0,0}$, noticing that $E[Y_{T}|D=d,T=t]=\lim_{h\rightarrow 0}E\left[\frac{\omega(D;d,h)\cdot I\{T=t\}\cdot Y_{T}}{\Pi_{d,t}}\right]$. For this reason, the DR expression (2) satisfies the Neyman orthogonality.
Next, consider the identification of the ATET $\Delta_{d_{t},d_{t}^{\prime},t}=E[Y_{t}(d_{t})-Y_{t}(d_{t}^{\prime})|D_{t}=d_{t}]$ in equation (7) based on Assumptions 4 to 7 using DR expression (2). The steps for showing Neyman orthogonality are equivalent when replacing the nuisance parameters in equation (A.1) by the following nuisance parameters, which additionally includes the treatment history $\mathbf{D_{t-1}}$ in the conditioning set:
$\eta=$$\{\mu_{d_{t}}(t-1,\mathbf{D_{t-1}},X),\mu_{d^{\prime}_{t}}(t,\mathbf{D_{t-1}},X),\mu_{d^{\prime}_{t}}(t-1,\mathbf{D_{t-1}},X),$
$\rho_{d_{t},t}(\mathbf{D_{t-1}},X),\rho_{d_{t},t-1}(\mathbf{D_{t-1}},X),\rho_{d^{\prime}_{t},t}(\mathbf{D_{t-1}},X),\hat{\rho}_{d^{\prime}_{t},t-1}(\mathbf{D_{t-1}},X)\}.$
In addition, consider the identification of the ATET $\Delta_{d_{t-s},d^{\prime}_{t-s},t}=E[Y_{t}(d_{t-s})-Y_{t}(d^{\prime}_{t-s})|D_{t-s}=d_{t-s}]$ in equation (11) based on Assumptions 4, 6, 7, and 8 using DR expression (2). Also in this case, the steps for showing Neyman orthogonality are analogous, with the only difference to the previous setup being the introduction of a time lag between the treatment and post-treatment outcomes.
Next, we demonstrate the Neyman orthogonality of the DR expressions (3) and (15) for panel data. Consider the identification of the ATET $\Delta_{d_{t},d_{t}^{\prime},t}=E[Y_{t}(d_{t})-Y_{t}(d_{t}^{\prime})|D_{t}=d_{t}]$ in equation (7) based on Assumptions 4 to 7. We define the nuisance parameters
$\eta=(m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}};X),p_{d_{t}}(X))$
and denote the true values of the nuisance parameters in the population by
$\eta^{0}=(m^{0}_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X),p^{0}_{d_{t}}(\mathbf{D_{t-1}},X)).$
We define $\Psi_{d^{\prime}}=E[m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)|D_{t}=d_{t}]$ to be the average of the conditional mean difference $m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)=E[Y_{t}-Y_{t-1}|D_{t}=d_{t}^{\prime},\mathbf{D_{t-1}},X]$ among those with treatment dose $D_{t}=d_{t}$. The ATET provided in equation (13) may be expressed as a function of $\Psi_{d^{\prime}}$:
$\Delta_{d_{t},d_{t}^{\prime},t}=E[m_{d_{t}}(t,\mathbf{D_{t-1}},X)-m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)|D_{t}=d_{t}]$
$=E[Y_{t}-Y_{t-1}|D_{t}=d_{t}]-\Psi_{d_{t}^{\prime}},$ (A.4)
where the second equality follows from the law of iterated expectations. Therefore, considering a Neyman-orthogonal function for $\Psi_{d_{t}^{\prime}}$ yields a Neyman-orthogonal expression of the ATET, which is an additive function of $\Psi_{d_{t}^{\prime}}$.
We denote by $\phi_{d_{t}^{\prime}}$ the Neyman-orthogonal score function for $\Psi_{d_{t}^{\prime}}$, which corresponds to the following expression:
$\phi_{d_{t}^{\prime}}$ $=\frac{\omega(D_{t};d_{t},h)\cdot m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)}{P_{d_{t}}}$
$+\frac{\omega(D_{t};d_{t}^{\prime},h)\cdot p_{d_{t}}(\mathbf{D_{t-1}},X)\cdot[Y_{t}-Y_{t-1}-m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)]}{p_{d^{\prime}_{t}}(\mathbf{D_{t-1}},X)\cdot P_{d_{t}}}-\Psi_{d_{t}^{\prime}}.$ (A.5)
The Gateaux derivative of $E[\phi_{d_{t}^{\prime}}]$ in the direction of $[\eta^{0}-\eta]$ is given by:
$\partial E[\phi_{d^{\prime},t^{\prime}}][\eta-\eta^{0}]=\lim_{h\rightarrow 0}E\left[\frac{\omega(D_{t};d_{t},h)\cdot[m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)-m^{0}_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)]}{P_{d_{t}}}\right]$
$-\lim_{h\rightarrow 0}E\left[\frac{\omega(D_{t};d_{t}^{\prime},h)\cdot p^{0}_{d_{t}}(\mathbf{D_{t-1}},X)\cdot[m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)-m^{0}_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)]}{p^{0}_{d^{\prime}_{t}}(\mathbf{D_{t-1}},X)\cdot P_{d_{t}}}\right]$ (A.6)
$+\lim_{h\rightarrow 0}E\left[\frac{\omega(D_{t};d_{t}^{\prime},h)\cdot[Y_{t}-Y_{t-1}-m^{0}_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)]}{p^{0}_{d^{\prime}_{t}}(\mathbf{D_{t-1}},X)}\cdot\frac{[p_{d_{t}}(\mathbf{D_{t-1}},X)-p^{0}_{d_{t}}(\mathbf{D_{t-1}},X)]}{P_{d_{t}}}\right]$
$-\lim_{h\rightarrow 0}E\left[\frac{\omega(D_{t};d_{t}^{\prime},h)\cdot p^{0}_{d_{t}}(\mathbf{D_{t-1}},X)\cdot[Y_{t}-Y_{t-1}-m^{0}_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)]}{p^{0}_{d^{\prime}_{t}}(\mathbf{D_{t-1}},X)\cdot P_{d_{t}}}\cdot\frac{p_{d^{\prime}_{t}}(\mathbf{D_{t-1}},X)-p^{0}_{d^{\prime}_{t}}(\mathbf{D_{t-1}},X)}{p^{0}_{d^{\prime}_{t}}(\mathbf{D_{t-1}},X)}\right]=0.$
Analogously to expression (A.1), applying the law of iterated expectations to the term in the second line of equation (A.1) yields:
$-\lim_{h\rightarrow 0}E\left[\frac{E[\omega(D_{t};d_{t}^{\prime},h)|\mathbf{D_{t-1}},X]\cdot p^{0}_{d_{t}}(\mathbf{D_{t-1}},X)\cdot[m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)-m^{0}_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)]}{p^{0}_{d^{\prime}_{t}}(\mathbf{D_{t-1}},X)\cdot P_{d_{t}}}\right]$
$=-E\left[\frac{p^{0}_{d_{t}}(\mathbf{D_{t-1}},X)\cdot[m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)-m^{0}_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)]}{P_{d_{t}}}\right].$
Applying the law of iterated expectations to the term on the left hand side of the first line yields:
$\lim_{h\rightarrow 0}E\left[\frac{E[\omega(D_{t};d_{t},h)|\mathbf{D_{t-1}},X]\cdot[m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)-m^{0}_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)]}{P_{d_{t}}}\right]$
$=E\left[\frac{p^{0}_{d_{t}}(\mathbf{D_{t-1}},X)\cdot[m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)-m^{0}_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)]}{P_{d_{t}}}\right].$
Consequently, the terms in the first and in the second line of equation (A.1) cancel out. Considering the third line, we have that:
$\lim_{h\rightarrow 0}E\left[E\left[\frac{\omega(D_{t};d_{t}^{\prime},h)\cdot[Y_{t}-Y_{t-1}-m^{0}_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)]}{p^{0}_{d^{\prime}_{t}}(\mathbf{D_{t-1}},X)}\cdot\frac{[p_{d_{t}}(\mathbf{D_{t-1}},X)-p^{0}_{d_{t}}(\mathbf{D_{t-1}},X)]}{P_{d_{t}}}\bigg{|}\mathbf{D_{t-1}},X\right]\right]$
$=[m^{0}_{d_{t}^{\prime}}(\mathbf{D_{t-1}},X)-m^{0}_{d_{t}^{\prime}}(\mathbf{D_{t-1}},X)]\cdot\frac{[p_{d_{t}}(\mathbf{D_{t-1}},X)-p^{0}_{d_{t}}(\mathbf{D_{t-1}},X)]}{P_{d_{t}}}=0.$
Concerning the fourth line,
-- $\lim_{h\rightarrow 0}E\left[E\left[\frac{\omega(D_{t};d_{t}^{\prime},h)\cdot p^{0}_{d_{t}}(\mathbf{D_{t-1}},X)\cdot[Y_{t}-Y_{t-1}-m^{0}_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)]}{p^{0}_{d^{\prime}_{t}}(\mathbf{D_{t-1}},X)\cdot P_{d_{t}}}\right.\right.$
×$\times\left.\left.\frac{p_{d^{\prime}_{t}}(\mathbf{D_{t-1}},X)-p^{0}_{d^{\prime}_{t}}(\mathbf{D_{t-1}},X)}{p^{0}_{d^{\prime}_{t}}(\mathbf{D_{t-1}},X)}\bigg{|}\mathbf{D_{t-1}},X\right]\right]$
$=$ $-E\left[\frac{p^{0}_{d_{t}}(\mathbf{D_{t-1}},X)\cdot[m^{0}_{d_{t}^{\prime}}(\mathbf{D_{t-1}},X)-m^{0}_{d_{t}^{\prime}}(\mathbf{D_{t-1}},X)]}{P_{d_{t}}}\cdot\frac{p_{d^{\prime}_{t}}(\mathbf{D_{t-1}},X)-p^{0}_{d^{\prime}_{t}}(\mathbf{D_{t-1}},X)}{p^{0}_{d^{\prime}_{t}}(\mathbf{D_{t-1}},X)}\right]=0.$
Thus, the Gateaux derivative of $E[\phi_{d^{\prime}_{t}}]$ in the direction of $[\eta^{0}-\eta]$ is zero. Furthermore, the Gateaux derivative of $E[Y_{t}-Y_{t-1}|D_{t}=d_{t}]$ in the direction of $[\eta^{0}-\eta]$ is zero. We note that the DR expression (3) is equivalent to expression (A.1) while using the Neyman-orthogonal score functions (A.1) for computing $\Psi_{d^{\prime}_{t}}$, noting that $E[Y_{t}-Y_{t-1}|D_{t}=d_{t}]=\lim_{h\rightarrow 0}E\left[\frac{\omega(D;d,h)\cdot(Y_{t}-Y_{t-1})}{P_{d_{t}}}\right]$. As a result, the DR expression (3) satisfies the Neyman orthogonality.
Lastly, consider the identification of the ATET $\Delta_{d_{t-s},d^{\prime}_{t-s},t}=E[Y_{t}(d_{t-s})-Y_{t}(d^{\prime}_{t-s})|D_{t-s}=d_{t-s}]$ in equation (11) based on Assumptions 4, 6, 7, and 8 using DR expression (15) for panel data. In this case, the steps for showing Neyman orthogonality are analogous. The only difference to the previous is the introduction of a time lag between the treatment and post-treatment outcomes.
A.2 Sketch of proofs of asymptotic theories
The discussion below sketches the main ideas of the proofs for our asymptotic results, which closely follow Zhang [2024]. We focus on the panel data case and note that the arguments extend to the repeated cross-sections case with only minor changes. We recall that under the regularity conditions of Section 5, we defined the bias as the difference between the true ATET and its kernel counterpart when the continuous treatment intensity is approximated by a kernel function with bandwidth $h$:
$B(h)=\Delta_{d_{t},d_{t}^{\prime},t}-\Delta_{d_{t},d_{t}^{\prime},t}^{h}.$
In the panel case, this bias can be shown to depend on the difference between the conditional density and its kernel approximation:
$\rho_{d}(D_{t-1},X)-\rho_{d}^{h}(D_{t-1},X).$
Moreover, we can demonstrate that
$\rho_{d}^{h}(D_{t-1},X)$ $=E[\omega(D;d,h)|\mathbf{D_{t-1}},X]$
[…]
$=\rho_{d}(D_{t-1},X)+Ch^{2}\rho_{d}^{(2)}(D_{t-1},X)+o(h^{2}).$
For this reason, taking the difference $\rho_{d}(D_{t-1},X)-\rho_{d}^{h}(D_{t-1},X)$ and making use of the assumptions on the kernel function and the conditional density, the bias is of order $O(h^{2})$.
Secondly, to show asymptotic normality, we may employ the following decomposition:
$\hat{\Delta}_{d_{t},d_{t}^{\prime},t}-\Delta_{d_{t},d_{t}^{\prime},t}$ $=\hat{\Delta}_{d_{t},d_{t}^{\prime},t}-\Delta_{d_{t},d_{t}^{\prime},t}^{h}+\Delta_{d_{t},d_{t}^{\prime},t}-\Delta_{d_{t},d_{t}^{\prime},t}^{h}$
$=\hat{\Delta}_{d_{t},d_{t}^{\prime},t}-\Delta_{d_{t},d_{t}^{\prime},t}^{h}+B(h).$
It suffices to focus on the difference between our ATET estimator and the kernel counterpart of the population ATET. By the definition of the score function (22) in Section 4, an application of Taylor’s theorem yields the following decomposition:
$\sqrt{n}(\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}-\Delta_{d_{t},d_{t}^{\prime},t}^{h})$ $=\sqrt{n}\sum_{i=1}^{n}E_{n}[\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\hat{\eta})]$ (A.7)
$+\sqrt{n}\sum_{i=1}^{n}E_{n}[\partial_{P}\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\hat{\eta})](\hat{P}_{d_{t}}-P_{d_{t}})$ (A.8)
$+\sqrt{n}\sum_{i=1}^{n}E_{n}[\partial_{P}^{2}\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\bar{P},\hat{\eta})](\hat{P}_{d_{t}}-P_{d_{t}})^{2}$ (A.9)
where the derivative $\partial_{P}$ is w.r.t. the density $P_{d_{t}}$, and $\bar{P}\in(\hat{P}_{d_{t}},P_{d_{t}})$. We use $E_{n}$ to denote the empirical average. Furthermore, we omit a discussion of cross-fitting since it does not affect our main argument. Then, the asymptotic result follows from standard arguments and we only provide a high-level summary. In the first step, we show that the second-order term (A.9) vanishes asymptotically and does not contribute to the asymptotic variance. In the second step, we bound the first-order term (A.8) and show that this term contributes to the asymptotic variance. In the third step, we expand (A.7) around the nuisance estimator $\hat{\eta}$ and use Neyman orthogonality to show that the first-order bias disappears and second-order terms are negligible. In the final step, we employ a triangular array central limit theorem to establish asymptotic normality.
To establish the consistency of our variance estimator, we define:
$\tilde{\phi}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}})=\phi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)-\frac{\Delta_{d_{t},d_{t}^{\prime},t}^{h}}{P_{d_{t}}}\left(\omega(D_{t};d_{t},h)-E[\omega(D_{t};d_{t},h)]\right).$
Note that by definition,
$\sigma_{h}^{2}=E[\tilde{\phi}^{2}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)];\quad\hat{\sigma}_{h}^{2}=E_{n}[\tilde{\phi}^{2}(W,\hat{\Delta}_{d_{t},d_{t}^{\prime},t},\hat{P}_{d_{t}},\hat{\eta})].$
It has to be shown that
$J_{n}=|E_{n}[\tilde{\phi}^{2}(W,\hat{\Delta}_{d_{t},d_{t}^{\prime},t},\hat{P}_{d_{t}},\hat{\eta})]-E[\tilde{\phi}^{2}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)]|=o_{p}(1).$
Using the triangle inequality, we have:
$J_{n}$ $\leq|E_{n}[\tilde{\phi}^{2}(W,\hat{\Delta}_{d_{t},d_{t}^{\prime},t},\hat{P}_{d_{t}},\hat{\eta})]-E_{n}[\tilde{\phi}^{2}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)]|$ (A.10)
$+|E_{n}[\tilde{\phi}^{2}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)]-E[\tilde{\phi}^{2}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)]|.$ (A.11)
We can demonstrate that (A.11) is $o_{p}(1)$ by making use of the assumptions on the kernel function and boundedness. Furthermore, one has to bound the term (A.10). We henceforth denote (A.10) by $J_{1n}$, which satisfies:
$J_{1n}^{2}\lesssim S_{n}\left(S_{n}+\frac{1}{n}\sum_{i=1}^{n}\tilde{\phi}^{2}(W_{i},\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)\right),$
where
$S_{n}=\frac{1}{n}\sum_{i=1}^{n}|\tilde{\phi}(W_{i},\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)-\tilde{\phi}(W,\hat{\Delta}_{d_{t},d_{t}^{\prime},t},\hat{P}_{d_{t}},\hat{\eta})|^{2}.$
Specifically, $S_{n}$ may be of order $O_{p}(h^{-1}\varepsilon_{n}^{2}+h^{-2}n^{-1})$, which follows from an expension of the term $\tilde{\phi}(W,\hat{\Delta}_{d_{t},d_{t}^{\prime},t},\hat{P}_{d_{t}},\hat{\eta})$ around $\hat{\Delta}_{d_{t},d_{t}^{\prime},t}$ using Taylor’s theorem. Moreover, the remaining term in $J_{1n}^{2}$, namely $\frac{1}{n}\sum_{i=1}^{n}\tilde{\phi}^{2}(W_{i},\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)$, may be of order $O_{p}(h^{-1})$ by our assumptions on the kernel function and boundedness. Therefore, $J_{1n}$ is $o_{p}(1)$ if $h^{-2}\varepsilon_{n}^{2}+h^{-3}n^{-1}=o(1)$, and the consistency of our variance estimator follows.